### v2

# Colorful Minors

Abstract
We introduce the notion of colorful minors, which generalizes the classical concept of rooted minors in graphs. A -colorful graph is defined as a pair where is a graph and assigns to each vertex a (possibly empty) subset of at most colors. The colorful minor relation enhances the classical minor relation by merging color sets at contracted edges and allowing the removal of colors from vertices. This framework naturally models algorithmic problems involving graphs with (possibly overlapping) annotated vertex sets. We develop a structural theory for colorful minors by establishing three core theorems characterizing -colorful minor-free graphs, where consists either of a clique or a grid with all vertices assigned all colors, or of grids with colors segregated and ordered on the outer face. Our results reveal that when exclusion is imposed not only on graphs but also to the way colors are distributed in them, a more refined structural landscape appears.
Leveraging our structural insights, we provide a complete classification – parameterized by the number of colors – of all colorful graphs that exhibit the Erdős–Pósa property with respect to colorful minors. On the algorithmic side, we deduce that colorful minor testing is fixed-parameter tractable. Together with the fact that the colorful minor relation forms a well-quasi-order, this implies that every colorful minor-monotone parameter on colorful graphs admits a fixed-parameter algorithm. Furthermore, we derive two algorithmic meta-theorems (AMTs) whose structural conditions are linked to extensions of treewidth and Hadwiger number on colorful graphs. Our results suggest how known AMTs can be extended to incorporate not only the structure of the input graph but also the way the colored vertices are distributed in it.
Keywords: Graph Minors, Colorful Minors, Annotated Graphs, Rooted Minors, Erdős-Pósa property, Structural Graph Theory, Obstruction sets, Algorithmic Meta-Theorems, Bidimensionality.
Contents
1 Introduction
Over the last decades, a prominent interplay between structural graph theory and theoretical computer science has been the search for structural properties of graphs that facilitate the design of efficient algorithms for otherwise NP-hard problems (see for example [18]). One of the most studied such parameters is a measure for the “tree-likeness” known in its most popular form as treewidth [75, 4, 5, 6]. Treewidth plays an exceptional role not only in the world of (parameterized) algorithms, but also in the theory of graph minors by Robertson and Seymour [73, 76, 77, 78]. A central theorem of Robertson and Seymour is the Grid Theorem [76] that states that any minor-closed graph class has bounded treewidth if and only if does not contain all planar graphs. This reveals a fundamental limitation in the use of treewidth (or any minor-monotone parameter really) as a parametrization for an NP-hard problem: While there exists a wide range of problems that are tractable on classes of bounded treewidth, there exist many important problems that are NP-hard even on planar graphs [34, 16]. This means that, in the regime of minor-closed graph classes, such problems are tractable if and only if a planar graph is excluded, unless .
A particularly interesting class of problems tractable on graphs of bounded treewidth, but NP-hard on more general minor-closed graph classes, are certain problems whose instances come with specific vertex sets, typically called roots, terminals, or annotated vertices. A prototypical example of such a problem is Steiner Tree. The Steiner Tree problem receives a graph accompanied with a special set of annotated vertices as input and asks for the minimum number of edges in a connected subgraph of containing all vertices of The problem is known to be NP-complete on planar graphs, however it can be solved in polynomial time with respect to structural parameterizations that instead of (or in addition to) restricting the global structure of the graph (like treewidth), focus on the structure relative to the annotation [22, 25, 55], thereby providing tractable algorithms beyond the regime of minor-closed classes of bounded treewidth.
This hints at the possibility of developing a general structural and algorithmic theory for annotated graphs, providing tools that supersede the limits imposed by classical graph minor theory.
Computational problems such as Steiner Tree can be naturally understood as problems on annotated graphs. Recently, several positive algorithmic results for Steiner Tree on annotated graphs excluding certain types of so-called rooted minors111An annotated graph is a rooted minor of an annotated graph if there exists a collection of pairwise vertex-disjoint connected subgraphs of such that for all there is an edge between and and for all we have have emerged [41, 46] alongside purely structural results on rooted minors [62, 44, 30] and found various applications [44, 2]. Moreover, Steiner Tree is just one example where a theory of rooted minors may open a new horizon of tractability. There exists a variety of other important problems exhibiting similar behavior, among them Multiway Cut [61, 19, 66], Disjoint paths [65, 82], #Defect Matching [17], and others [33, 16, 58]. Moreover, annotation has recently be shown to also play a key role in structural extensions of Courcelle’s Theorem [15, 13, 14] by restricting quantification to sets of low bidimensionality222The bidimensionality of a vertex set in a graph , denoted by , is the largest such that contains the -grid with all vertices annotated as a rooted minor [84]. [81].
Another area where annotation comes into play is the study of Erdős-Pósa dualities, which lies at the crossroads of structural and algorithmic graph theory. The Erdős-Pósa property for (rooted) minors asks, for a fixed (annotated) graph , whether for any (annotated) graph either contains many vertex-disjoint occurrences of as a (rooted) minor, or has a small vertex set whose deletion removes all occurrences of as a (rooted) minor. Note that not every (annotated) graph satisfies the Erdős-Pósa property. The Erdős-Pósa property has been investigated by many authors in a variety of settings [76, 68], and several key cases are also known for rooted minors [49, 70, 59, 44].
While the concept of rooted minors in annotated graphs appears to be relatively well explored, a general theory for deriving results such as those above still seems to be lacking. Beyond the fact that many of the results mentioned are inspired by the theory of graph minors, no systematic extension of its powerful toolkit has yet emerged. This is somewhat surprising given the significant structural depth and algorithmic potential of these concepts, as we elaborate below.
1.1 Our results
In this paper, we develop a comprehensive theory for a generalization of rooted minors in multi-annotated graphs, which we refer to as colorful minors of colorful graphs. Instead of a single annotated set of vertices we consider sets of annotated vertices, which may be overlapping. We investigate both the algorithmic and structural properties of colorful graphs that exclude certain classes of colorful minors – an appropriate adaptation of the concept of rooted minors to our setting. Our results suggest that much of the combinatorial and algorithmic power of classical graph minor theory can be effectively extended to the colorful setting.
From structure… Our core contributions are a series of structural theorems for -colorful minor-free -colorful graphs, both for general and for the particular case where is planar – under two distinct assumptions on the distribution of colors. Our results use as a departure point the notion of the bidimensionality of a vertex set and related structural concepts [68, 67, 69]. Building on our structural theorems, we derive a complete classification of all -colorful graphs that satisfy the Erdős-Pósa property, for every (under the colorful minor relation).
…to algorithms. Our next step is to apply our structural theorems to algorithmic settings, with the broader goal of developing an algorithmic graph structure theory applicable to problems involving vertex annotations. To this end, we first observe that colorful graphs are well quasi-ordered under the colorful minor relation. We also provide an algorithm that decides whether a -colorful graph is a colorful minor of a -colorful graph in time333By we mean that there is a function such that Also, given a graph we use and for the number of its vertices and edges respectively. These two facts imply that membership for any colorful minor-closed class of colorful graphs is decidable in polynomial time.
Next, we show how classical algorithmic meta-theorems (AMTs) – such as Courcelle’s Theorem and the recent algorithmic meta-theorem of Sau, Stamoulis, and Thilikos [81] – can be extended to colorful graphs, where the structural conditions are expressed via the exclusion of colorful minors. This framework enables the resolution of problems on graphs containing arbitrarily large grid or clique minors while, up to now, the exclusion of a minor appeared as a necessary (sparsity) condition.
Our results highlight a key insight:
Problem complexity does not only depend on the global structure of the input graph,
but crucially on how the annotated sets are distributed within it.
The algorithmic results presented here are not intended to be exhaustive; rather, they serve to illustrate this principle and to lay the groundwork for future exploration of algorithmic problems involving vertex annotations. In particular, we provide algorithmic applications for two of our three main structural results – Theorems˜1.2 and 1.3 – while leaving the investigation of algorithmic consequences of theorem˜1.4 to future work.
Colorful graphs.
Let be a non-negative integer. A -colorful graph is a pair where is a graph and444For we denote by and by For a vertex we call the palette of For illustrations, we use magenta for blue for , yellow for and green for
Certainly, every vertex of receives a (possibly empty) subset of colors. We denote the set of colors assigned to the vertices of by (for a vertex set we write for the set ). Similarly, for we denote by the set When is a singleton, we write instead of
A -colorful graph where is called restricted and a -colorful graph with is called empty. Note that any restricted -colorful graph is also a -colorful graph, so the notion of being restricted is highly dependent on the context given by the total set of colors. Finally, a -colorful graph is said to be rainbow555Notice that – in slight violation of the English language – we use the word “rainbow” as an adjective here. if for all
A colorful graph is a (colorful) subgraph of a colorful graph denoted by if is a subgraph of and for all If is a colorful graph and is a subgraph of we write, in slight abuse of notation, for the -colorful subgraph of where all vertices of receive precisely the colors they received in
A colorful graph is a colorful minor of a colorful graph if can be obtained from by means of the following operations:
deleting an edge
deleting a vertex and restricting the domain of to
for some and overwrite (we refer to this operation as removing a color (from )), and
contracting an edge that is introducing a new vertex with neighborhood setting and then deleting both and from the resulting colorful graph.
See figure˜1 for an example. Notice that -colorful graphs are exactly annotated graphs and the notions of colorful minors and rooted minors coincide on annotated graphs.
In section˜2.4 we derive the following, as a simple consequence of a deep theorem of Robertson and Seymour [74].
Theorem 1.1.
For every non-negative integer the class of all -colorful graphs is well-quasi-ordered by the colorful minor relation.
Structural results.
We prove a total of three structure theorems as follows. The first two describe the structure of graphs excluding the -colorful rainbow -clique and the -colorful rainbow -grid respectively as a colorful minor. Moreover, we generalize a theorem of Marx, Seymour, and Wollan [62], that describes the structure of -colorful graphs excluding a -grid where exactly the vertices of the first row receive color to an arbitrary number of colors.
We begin with our simplest result – excluding a rainbow clique – which can be naturally phrased in terms of a modulator to restricted graphs. For a vertex set in a graph we define the torso of in denoted by as the graph obtained from by turning the neighborhood of in into a clique for every component of In short, every -colorful graph that excludes a rainbow clique as a colorful minor has a set whose torso excludes a minor and where every component of is free of some (not necessarily the same) color.
Theorem 1.2.
There exists a function such that for all non-negative integers and and all -colorful graphs one of the following holds:
i)
contains a rainbow -clique as a colorful minor, or
ii)
there exists a set such that the torso of in is -minor-free and for all components of the -colorful graph is restricted.
Moreover, and there exists an algorithm that finds one of the two outcomes above in time
A tree-decomposition for a graph is a pair where is a tree and called the bags of assigns a set of vertices of to every node of such that for every there is with and for every the set is connected. The adhesion of is and the width of is defined as The treewidth of a graph denoted by is the smallest integer such that there is a tree-decomposition of width at most for
Our next theorem generalizes a theorem of Marx, Seymour, and Wollan [62] (see also [44]). Let be an integer. A -segregated grid is a -colorful graph where is the -grid where the vertices of the first column can be numbered as in order of their appearance, and we have for all and there exists a permutation of such that for every In this situation, we say that realizes . See figure˜2 for an example.
Theorem˜1.2 explains that every -colorful minor-free colorful graph can be turned into a collection of restricted graphs by deleting an -minor-free subgraph. Segregated grids capture exactly the situation where we require to be a planar graph. We may express this in form of a parameter as follows. The torso treewidth of a -colorful graph is the smallest integer such that there exists a set such that the treewidth of the torso of in is at most and for every component of the -colorful graph is restricted.
Theorem 1.3.
There exists a function such that for all non-negative integers and and all -colorful graphs one of the following holds:
i)
contains as a colorful minor some -segregated grid and the torso treewidth of is more than , or
ii)
the torso treewidth of is at most
Moreover, and there exists an algorithm that finds one of the two outcomes above in time
Notice that theorem˜1.3 is optimal in the sense that the torso treewidth of segregated grids is unbounded (see lemma˜5.4). In other words, theorem˜1.3 can be seen as a mix-max duality theorem for segregated grids. We remark that for the case the torso treewidth of has been independently defined by Jansen and Swennenhuis [46] and by Hodor, La, Micek, and Rambaud [45] under different names.
It is easy to see that theorem˜1.3 is far from a structure theorem for -colorful graphs that exclude an arbitrary planar colorful minor. This is in stark contrast to the -colorful case provided by the Grid Theorem [76]. It turns out that, in order to fully describe the structure of -colorful graphs that exclude a fixed but arbitrary planar colorful minor, we need the full power of Robertson and Seymour’s Graph Minor Structure Theorem (GMST) [72, 53, 38]. The definitions necessary to state the GMST are rather involved and technical, for an in-depth explanation see section˜4.
We say that a graph has a -near embedding in a surface if there exists with such that such that has an embedding into with pairwise vertex-disjoint faces such that for all the ’s are pairwise vertex-disjoint for positive and for each has a path-decomposition 666A path-decomposition of a graph is a tree-decomposition where is a path. of width at most such that the vertices of appear in agreement to their cyclic ordering on the boundary of and for all The set is called the apex set, the graphs are called the vortices, and the sets are the interiors of the vortices.
In the context of -colorful graphs, we need some additional information regarding the colors. Let be a tree decomposition of a -colorful graph The colorful torso of at a node is the colorful graph obtained from by turning, for every the set into a clique and adding to for all where is the unique component of that contains
Theorem 1.4.
There exists a function such that for all non-negative integers and and all -colorful graphs one of the following holds:
i)
contains the -colorful rainbow -grid as a colorful minor, or
ii)
has a tree-decomposition of adhesion at most such that for all either is a leaf with unique neighbor and is restricted, or the colorful torso of at has an -near embedding and there exists such that all vertices with belong to the apex set or the interior of a vortex.
Moreover, and there exists an algorithm that finds one of the two outcomes above in time
Notice that theorem˜1.4 can also be seen as a min-max duality as implied by results of Thilikos and Wiederrecht [84].
To rephrase theorem˜1.4, the absence of a rainbow grid as a colorful minor means that, with respect to any large grid-minor in and in the absence of a big clique minor – where a local version of theorem˜3.6 would hold – there is always at least one color of small bidimensionality and all vertices of such colors may be confined to the apex set and the vortices of a near-embedding. This means that for -colorful graphs, all colored vertices are fully confined to the vortices and the apex set and the entire surface part of the near embedding does not have any colored vertices. Indeed, this appears to be a far-reaching generalization of the idea to cover all colored vertices in a planar graph with few faces. We believe this provides a theoretical foundation on why many problems that are tractable on planar graphs when their annotated set is covered by few faces, remain tractable on all colorful minor-closed classes of -colorful graphs that exclude some fixed planar -colorful graph. It moreover indicates that, in algorithm design, vortices in -near embeddings may play a similar role to the one played by faces in planar graphs.
Delineating the Erdős-Pósa property for colorful minors.
We say that a -colorful graph has the Erdős-Pósa property if there exists a function such that for every every -colorful graph either has pairwise vertex-disjoint -colorful subgraphs, each containing as a colorful minor, or there is a set with such that does not contain as a colorful minor.
It follows directly from a construction by Robertson and Seymour [76] that if a pair has the Erdős-Pósa property, then the graph must be planar. As we demonstrate in section˜6, the boundary between instances that do and do not exhibit the Erdős-Pósa property shifts in subtle ways in the presence of colors – subtleties that are worth highlighting.
For every we provide the full list of graphs such that a -colorful graph has the Erdős-Pósa property if and only if it does not contain any of the graphs from as a colorful minor. In other words, we provide, for every the set of colorful minor-obstructions for the Erdős-Pósa property of -colorful graphs. We also stress that i.e., its size grows polynomially in the number of colors. Figure˜3 gives a schematic definition of for every
Theorem 1.5.
For every non-negative integer a -colorful graph has the Erdős-Pósa property if and only if it does not contain any of the graphs in as a colorful minor.
theorem˜1.5 illustrates that the excluded minors for -colorful graphs do not need to have all colors. In fact contains graphs without any colored vertices and no member of has five colors no matter the choice of Our results on the Erdős-Pósa property indicate that colorful minors are a proper generalization of the concept of minors, able to express a variety of different properties.
Towards a general framework for colorful parameters.
With theorem˜1.3 and theorem˜1.4 in hand, it becomes apparent that there are fundamentally distinct ways to extend a graph parameter to a colorful analogue, and that these choices can lead to remarkably different structural decompositions. Which extension should be regarded as “canonical” appears to depend strongly on the intended application and on the structural/algorithmic consequences one wishes to capture.
For instance, from the perspective of structural decompositions, torso treewidth emerges as a particularly natural analogue of treewidth, and this is further supported by the fact that it delineates the Erdős–Pósa property for colorful minors as theorem˜1.5 shows, mirroring the role played by treewidth in the uncolored setting. In contrast, when viewed through the lens of obstructions, bidimensionality provides the appropriate analogue, as it yields the correct structure theorem for colorful graphs excluding an arbitrary colorful planar graph as a colorful minor, and potentially, the correct structural parameterization to obtain precise dichotomy complexity results for problems on colorful graphs that become NP-hard in planar graphs, as previously discussed.
These observations point towards a broader, open-ended research direction, which is to better understand these phenomena by examining other classical graph parameters beyond treewidth, and by developing general principles for lifting such parameters to colorful analogues. We return to this perspective in section˜8.2 of the conclusion, where we discuss this issue in more detail and propose a conjecture.
Algorithmic results.
One of the ground-breaking results of Robertson and Seymour’s Graph Minors Series is the Graph Minor Algorithm that allows to check if a graph is a minor of a graph in time [79] (see [54, 56] for improvements). This algorithm also solves a rooted version of the Minor Checking problem which, as a special case, contains the so-called -Disjoint Paths problem.
The -Colorful Minor Checking problem has as input two -colorful graphs and and the task is to determine whether contains as a colorful minor.
We present a simple reduction to the rooted minor checking problem which allows us to leverage the algorithm of Korhonen, Pilipczuk, and Stamoulis [56] in order to deduce a similar result for -Colorful Minor Checking.
Theorem 1.6.
There exist an algorithm that takes as input two -colorful graphs and and solves -Colorful Minor Checking in time
Theorems˜1.1 and 1.6 imply that the recognition problem for any proper colorful minor-closed class can be solved in polynomial time. We present below a parametric variant of this, based on the general framework introduced by Fellows and Langston [27, 28] for the existence of fixed-parameter tractable algorithms.
A (-colorful) graph parameter is a map assigning a non-negative integer to every (-colorful) graph. We say that a -colorful graph parameter is colorful minor-monotone if for every -colorful graph and every colorful minor of we have Again, combining Theorems˜1.1 and 1.6 we can deduce the following.
Theorem 1.7.
For every colorful minor-monotone -colorful graph parameter there exist an algorithm that, given a -colorful graph and a non-negative integer checks whether in time.
We stress that the result of theorem˜1.7 is not constructive, i.e., it does not give any way to construct the claimed algorithm, as is is based on the finiteness of the obstruction set of the class which is only existentially provided by theorem˜1.1.
Meta-algorithmic Applications.
Our two algorithmic results can be seen as extensions of Courcelle’s Theorem and of a recent result due to Sau, Stamoulis, and Thilikos [81] to colorful graphs. While the structural parameters in these theorems are treewidth, denoted by and Hadwiger number, denoted by (that is the maximum size of a clique minor) respectively, we now introduce two parameters on colorful graphs that serve as our new combinatorial restrictions.
A key condition for applying our algorithmic results is the notion of folio-representability of a class of colorful graphs (i.e., the yes-instances of some algorithmic problem), formally defined in section˜7. Informally, is folio-representable if one can efficiently replace a subgraph of a colorful graph with colored vertices only within a small subset of the vertices of – viewed as a “protuberance”777We use the term protuberance to distinguish it from the term “protrusion” introduced in [7], where similar replacement operations were applied. Unlike protrusions, which are required to have bounded treewidth, protuberances are characterized by being “internally colorless”. interacting with the rest of the graph via the “boundary” – by a bounded-size subgraph such that the resulting graph belongs to if and only if does.
This replacement is guided by a notion of folio-equivalence: and must have the same set of rooted topological minors (up to a given size), with respect to the specified boundary Efficient computation of such replacements is enabled by results from [42], making folio-representability a natural combinatorial condition for the class
In what follows, whenever we refer to 1-colorful graphs we may alternatively use the notation instead of where . Given a graph and a set , we define the monodimensionality of in as the maximum order of a -segregated grid colorful minor of Given a -colorful graph and a color we define the monodimensionality of in as the monodimensionality of in
We now give the logical conditions under which our framework operates. In our first result, we require the class to be definable in Counting Monadic Second-Order Logic888There exists a formula such that if and only if models (CMSO). Courcelle’s Theorem says that, for every graph class definable by some CMSO-formula membership in can be decided in time We prove the following analogue for colorful graphs.
Theorem 1.8.
Let and let be a -colorful graph problem that is both folio-representable and definable by some CMSO-formula Then, there exists an algorithm that, given a -colorful graph decides whether in time where is the maximum monodimensionality of a color in .
From theorem˜1.3, the maximum monodimensionality of a color in a -colorful graph is equivalent to the maximum torso treewidth of over all colors Notice that theorem˜1.8 can be restated by setting to be this equivalent quantity, which plays the same role for colorful graphs as treewidth does for graphs in Courcelle’s Theorem.
For our second result, we consider a fragment of denoted as In this logic, quantification is restricted to sets of bounded bidimensionality, and a special predicate is available, asserting the existence of vertex-disjoint paths connecting the pairs It was shown by Sau, Stamoulis, and Thilikos [81] that if a graph class is definable in then membership in can be decided in time .
The Hadwiger number of a color in a -colorful graph is the maximum for which contains the rainbow -clique as a colorful minor. We prove the following.
Theorem 1.9.
Let and let be a -colorful graph problem that is folio-representable and definable by some -formula Then, there exists an algorithm that, given a -colorful graph decides whether in time , where is the maximum Hadwiger number of a color of .
We emphasize that the applicability of Theorems˜1.8 and 1.9, based on the structural theorems Theorems˜1.2 and 1.3 respectively, exceeds the scope of Courcelle’s Theorem and the results of [81]. The generality of our extension is displayed by the fact that, for every , there are classes of -colorful graphs for which the maximum monodimensionality/Hadwiger number of a color in is bounded while the treewidth/Hadwiger number of is unbounded.
This is a strong indication that, for problems on colorful graphs, the appropriate structural parameterizations toward deriving (meta-)algorithmic results should take into account the way the colors are distributed in the inputs.
We derive Theorems˜1.8 and 1.9 in section˜7 as special cases of a more general result applicable to optimization problems (theorem˜7.4). In fact, in section˜7, we see these theorems as corollaries of broader statements (theorem˜7.9 and theorem˜7.10) formulated in terms of optimization problems where we additionally display the parameter dependencies hidden in the -notation.
1.2 Brief insight into our approach
The general approach to proving Robertson-Seymour-style structure theorems typically constists of establishing a local structure theorem with respect to a wall and then applying a well-known technique, originating from the work of Robertson and Seymour [78], that turns the local statement into a global one based on tree-decompositions. Our structure theorems 1.2, 1.3, and 1.4 follow this general paradigm.
In the uncolored setting, proving a local structure theorem with respect to a wall usually proceeds as follows: Starting from a large wall in the graph , one either finds
a.i)
a large clique-minor that is highly connected to or
a.ii)
a structural decomposition of relative to typically of a topological nature, such as a flat wall or, more generally, a near embedding of in some surface with additional properties tailored to the theorem at hand.
For our results, handling outcome (a.ii) relies on sophisticated topological arguments stemming from the Graph Minor Structure Theorem [72] and its subsequent improvements [53, 38] as well as more recent refinements [83, 68, 69]. While technically demanding, this part largely builds on and adapts existing machinery to deal with annotation in near embeddings [69], except for theorem˜1.3 which uses different techniques which we comment on in the next paragraph.
Rainbow clique-minors and a multicolored Menger-type theorem.
The genuinely new difficulties and contributions arise in outcome (a.i). In the colored settings, the mere presence of a large clique-minor provides little information on the structure of the colorful graph with respect to the colored vertices. For our purposes, it is therefore essential to understand how well the colored vertices are connected to this clique-minor. Concretely, given a -colorful graph containing a large clique-minor, we aim to establish the following dichotomy:
b.i)
there exists a small set of vertices and a color such that the component of containing most of the clique-minor is free of vertices of color or
b.ii)
contains a large rainbow clique-minor that is highly connected to the original clique-minor.
This results, formalized in theorem˜3.1, is proved in section˜3.1 and constitutes a central new ingredient underpinning all our structural theorems.
In the context of the Disjoint Paths problem, a similar result was established by Robertson and Seymour [79] for -colorful graphs with a bounded number of colored vertices. Our setting is substantially more general: we allow -colorful graphs with an arbitrary number of colored vertices. Our proof involves a one-to-many strengthening of Menger’s Theorem, stated as lemma˜3.2, which may be of independent interest. For all positive integers , given vertex sets and in a graph either
c.i)
there exist disjoint paths in with exactly paths linking to or
c.ii)
there exists an index and a vertex set of size less than that separates from
Interestingly, our results on the Erdős-Pósa property for colorful minors imply that even a two-to-two strengthening of Menger’s theorem cannot hold in general, not even approximately (see section˜8.1). Consequently, our structural results are in a sense “tight”, as we cannot expect to replace condition (b.i) with a stronger alternative.
It is noteworthy that in the local version, theorem˜5.2, of theorem˜1.3, outcome (a.ii) also relies on the one-to-many strengthening of Menger above to control how well-connected the colored vertices are to a given flat wall in the graph.
Annotations and the Erdős-Pósa property.
Erdős-Pósa-type dualities in graph minors and beyond are intrinsically related to structural characterization of restricted classes of graphs, often in terms of “global” decomposition theorems based on tree-decompositions. A prime example of this interaction is the celebrated characterization of the proper minor-closed classes which have the Erdős-Pósa property for minors, shown by Robertson and Seymour [76] to be precisely those of bounded treewidth.
Crucial to this duality is showing that graph classes of unbounded treewidth necessarily fail to satisfy the Erdős–Pósa property. By the grid theorem, such classes are precisely those containing all planar graphs. Consequently, this reduces to understanding why the two Kuratowski minors and themselves do not admit the Erdős–Pósa property. In [76], Robertson and Seymour address this by constructing families of graphs based on and that explicitly violate the Erdős–Pósa property. These constructions are inherently topological: The resulting graphs embed minimally in the torus or the projective plane, but not in the sphere. At a high level, the argument exploits the additivity of Euler-genus under disjoint union [63], which implies that repeatedly combining non-planar graphs forces unbounded Euler-genus.
In the colorful setting, we initially follow the same high-level strategy. We identify torso treewidth as the correct parameter governing the Erdős–Pósa property for colorful minors. Utilizing our grid theorem for torso treewidth (theorem˜1.3), we identify the set of -colorful graphs and for the reverse direction of theorem˜1.5, we have to show that each -colorful graph in violates the Erdős–Pósa property for colorful minors.
At this point, however, the classical approach breaks down. Unlike the uncolored case, most graphs in are planar with particular assumptions on the topology induced by the colored vertices. As a consequence, the standard arguments are no longer applicable, as they fundamentally rely on non-planar embeddings.
Instead, our arguments crucially exploit the distribution of colored vertices within these graphs. We show that the topology induced by this distribution—rather than by the underlying graph alone—can still be leveraged to construct families that violate the Erdős–Pósa property. Once more, this highlights the importance of the arrangement of colored vertices within the underlying graph.
1.3 Related work and organization of the paper
Related work.
Various combinatorial problems concerning rooted minors with specific root configurations have been studied by Kawarabayashi [50], Wollan [91], Fabila-Monroy and Wood [26], and Moore [64]. For -colorful graphs, structural results have been obtained by Marx, Seymour, and Wollan [62]; Hodor, La, Micek, and Rambaud [44]; and Fiorini, Kober, Seweryn, Shantanam, and Yuditsky [30].
Building on the structural result of Marx, Seymour, and Wollan [62], Hodor, La, Micek, and Rambaud [44] established the Erdős–Pósa property for all connected planar -colorful graphs in which the colored vertices lie on the boundary of a single face. For -colorful graphs, the only known result concerning the Erdős–Pósa property involves cycles spanning a prescribed number of distinct colors, as shown by Kwon and Marx [59]. In the more intricate setting of labeled minors, Erdős–Pósa dualities have also been established by Bruhn, Joos, and Schaudt [10].
The most extensively studied problem on -colorful graphs is the Steiner Tree problem, which was shown to be solvable in polynomial time when the torso treewidth is bounded [46], and when a rainbow is excluded [41]. To the best of our knowledge, no combinatorial or algorithmic study has been undertaken on -colorful graphs as a general and unifying combinatorial framework.
Organization of the paper.
section˜2 introduces all preliminary concepts we require throughout the paper. section˜2.3 particularly introduces key concepts from from graph minor theory for handling the structure of -minor-free graphs. Moreover, section˜2.4 establishes our well-quasi-ordering result and section˜2.5 the reduction for colorful minor checking. Afterwards, section˜3, section˜4, and section˜5 concern our structural results. section˜3 deals with the proof of the full version of theorem˜1.2, namely theorem˜3.6, on the structure of colorful graphs excluding a rainbow clique as a colorful minor. section˜4 proceeds with the full version of theorem˜1.4, namely theorem˜4.3, on the structure of colorful graphs excluding a rainbow grid as a colorful minor. The last section on structure results, section˜5, deals with the structure of graphs excluding segregated grids, in the form of theorem˜5.3, which is the full version of theorem˜1.3. section˜6 concerns our results on the Erdős-Pósa property for colorful graphs. section˜6.3 deals with the connected case of the positive part of theorem˜1.5, while in section˜6.4 we lift this result to the non-connected case. Finally in section˜6.5, we complete this section with the negative part of theorem˜1.5. Going into section˜7, we discuss algorithmic applications of our colorful minor framework and we present our two AMTs, namely theorem˜7.9 and theorem˜7.10, the full versions of theorem˜1.8 and theorem˜1.9 respectively. In section˜8 we discuss possible directions for future research. Last but not least, in Appendix A we present a (non-exhaustive) list of problems in which our AMTs apply.
2 Preliminaries
By we denote the set of integers and by the set of reals. Given any two integers we write for the set Notice that the set is empty whenever For any positive integer we set
Separations and linkages.
Let be a graph. A separation of is a tuple such that and there is no edge with one end in and the other in A set is called a separator of if has strictly more components than Given sets a set is an --separator if no component of contains both a vertex of and a vertex of Note that, somewhat counterintuitively, this means that an --separator in is not necessarily a separator of
Let be a graph and A path is said to be an --path if the endpoints of are and Let A path is said to be an --path if there exist and such that is an --path and no internal vertex of belongs to An --linkage is a collection of pairwise vertex-disjoint --paths. A linkage is a collection of paths such that there exist for which is an --linkage. The order of a linkage is the number of paths in In a slight abuse of notation, we identify a linkage and the graph
Proposition 2.1 (Menger’s Theorem).
For every positive integer every graph and every choice of sets there either exists an --linkage of order or there exists an --separator of size less than in
Moreover, there exists an algorithm that finds, given and either a linkage of order or an --separator of size less than in time
2.1 Minors
Let and be graphs. We say that is a minor of if can be obtained from by a sequence of (i) vertex deletions, (ii) edge deletions, and (iii) edge contractions.
It is easy to see that is a minor of if and only if there exists a collection of connected and pairwise vertex-disjoint subgraphs of called the branch sets, such that for every there exists an edge with and We call a minor model of in
Now let and be colorful graphs. We say that a minor model of in is a colorful minor model of in if for every and every there is some vertex such that We often drop the specifier “colorful” and just say that is a minor model of in
2.2 Highly linked sets and tangles
Tangles.
Let be a graph and be an integer. We denote by the set of all separations of order less than in An orientation of is a set such that for every exactly one of and belongs to
A tangle of order in a graph is an orientation of such that for all it holds that
Let and be tangles in We say that is a truncation of if
Let be a tangle of order in a graph and let be a set of size at most Notice that every separation of order at most in corresponds to a separation of order at most in It follows that is a tangle of order in Now suppose that has components Then there exists such that for all separations of order in We call the -big component of
Walls, cliques, and tangles.
Let be a positive integer and be a minor model of in Notice that for every separation there exists a unique such that for some We call the -big side of Moreover, if we let be the orientation of obtained by taking all separations such that is the -big side of then is a tangle. We call the tangle induced by and denote it by
Let be two positive integers. The -grid is the graph with the vertex set and the edges
The elementary -wall is in turn derived from the -grid by deleting all edges in the following set
and removing all vertices of degree at most 1 in the resulting graph. An -wall is a subdivision of the elementary -wall. An -wall is an -wall and a wall is any graph that is an -wall for some Notice that every wall with at least two columns and rows is -connected.
Notice that any -wall consists of pairwise-disjoint paths and pairwise-disjoint paths such that is a non-empty path for all each meets the s in order, and each meets the s in order. We call the the horizontal paths of and the the vertical paths of (sometimes we also refer to them as the rows and columns of respectively). The perimeter of is the unique cycle contained in and every vertex of not contained in is said to be an inner vertex of
A brick of a wall is any facial cycle of that consists of exactly six degree- vertices of
Let with let be a graph, and be an -wall in Let be the orientation of such that for every the set contains the vertex set of both a horizontal and a vertical path of we call the -majority side of Then is the tangle induced by If is a tangle in we say that controls the wall if is a truncation of
Let and be (colorful) graphs as well as be a tangle in We say that a minor model of in is controlled by if there does not exist a separation of order less than and an such that
-linked sets.
Let Moreover, let be a graph and be a vertex set. A set is said to be an -balanced separator for if for every component of it holds that Let be a non-negative integer. We say that is a -linked set of if there is no -balanced separator of size at most for in
Given a -linked set of we define
Then is a tangle of order in
Finally, we need an algorithmic way to identify highly linked sets, and find, given a highly linked set, a large wall whose tangle is a truncation of the tangle induced by the highly linked set. The first is a classic result due to Reed [71]. The second is done in [85] by turning to an algorithm a proof of Kawarabayashi, Thomas, and Wollan, from [53].
Proposition 2.2 (Reed [71]).
There exists an algorithm that takes as input an integer a graph and a set of size at most and finds, in time , either a -balanced separator of size at most for or correctly determines that is -linked in
Proposition 2.3 (Thilikos and Wiederrecht [85] (see Theorem 4.2.)).
Let be an integer and There exist universal constants and an algorithm that, given a graph and a -linked set computes in time a -wall such that is a truncation of
2.3 Graph minor structure
We continue by briefly introducing some key concepts for handling the general structure of -minor-free graphs. The definitions we introduce here act as a strict simplification of the notions and terminology used in most proofs of the Graph Minor Structure Theorem by Robertson and Seymour. For more in-depth discussions on this topic, the reader is referred to [72, 53, 38]. For the purposes of this paper, the simplified definitions are fully sufficient as we are mostly using the corresponding theorems as black boxes. Most of the notions below are adapted from [52].
By a surface we mean a compact -dimensional manifold with or without boundary.
Given a surface and a set we denote by the closure of and by we denote the set
A painting in a surface is a pair where is finite, has finitely many arcwise-connected components, those are called the cells of and for every cell the closure is a closed disk where If we call a vortex. we define and denote the set of cells of by
Notice that the cells of a painting define a hypergraph whose vertices are precisely the elements of called the nodes, and a node belongs to some cell if and only if
-renditions.
Let be a graph and be a surface. A -rendition of is a triple where
is a painting in
assigns to each cell a subgraph of and
is an injection
such that
- (R1)
- (R2)
for all distinct and are edge-disjoint,
- (R3)
for every cell and
- (R4)
for every cell
We write for the set and similarly, we lift the set of cells from to If the -rendition of is understood from the context, we usually identify the sets and along for ease of notation.
Vortices, linear decompositions, and depth.
Let be a graph, be a surface and be a -rendition of Moreover, let be a vortex.
Let be the vertices of in numbered corresponding to the order of encountering them when traversing along in anti-clockwise (or clockwise) direction starting from A linear decomposition of is a sequence of sets such that
- (V1)
and for all
- (V2)
and for all there exists such that and
- (V3)
for every the set forms an interval in
The adhesion of a linear decomposition is The width of a linear decomposition is The depth of a vortex is the smallest integer such that has a linear decomposition of adhesion at most and the width of is the smallest integer such that has a linear decomposition of width at most
The breadth of a -rendition of a graph is the number of vortices of the depth of is the maximum depth over all vortices of and finally, the width of is the maximum width over all of its vortices.
Grounded subgraphs, traces, and flat walls.
Let be a surface and be a graph with a -rendition The simple torso of under is the multi graph obtained from the subgraph by adding an edge between any pair of vertices for every non-vortex cell such that
Notice that is a graph that can be embedded in such that for every with the vertices of form a triangle bounding a face as follows: For any pair of adjacent vertices there exists at least one cell such that From here on let be any such cell. Moreover, there exists a curve between that is completely contained in and does not intersect any other element of In case is unique. we apply the same construction for cells with here, there are two possible choices. Here, we will always implicitly assume that comes with a unique selection for each such curve
Let be a cycle or path. The image of in is the graph such that
and
if and only if there exists a subpath of between and that does not contain any other vertex of
Notice that any subpath between and in such as the one from the second item of the above definition must be fully contained in a single cell of Hence, for any subpath of between vertices of that contains no other element of there exists a unique curve mirroring We may identify the edges of and the curves mirroring the subpaths of corresponding to those edges. For every edge we say that the curve mirroring is the drawing of
If is a path, we say that is grounded if has at least one edge. In case is a cycle we say that is grounded if is a simple999That is, if has at least three vertices and no parallel edges. cycle. If is grounded, we call the subset of that is the union of together with the union of the drawings of all edges of the trace of Notice that the trace of is a simple curve in case is a path and a closed curve in case is a cycle.
A -connected subgraph of is said to be grounded if every cycle of is grounded.
As before let be a surface and be a graph with a -rendition Now let be an -wall, in such that is grounded in We say that is flat in if there exists a disk such that intersects only in no vortex of is contained in all vertices of are contained in and the trace of the perimeter of separates all inner vertices of from
We say that a wall is flat in if there exists a -rendition of for some non-empty surface such that is flat in
The following theorem is known as the Flat Wall Theorem. It first appeared in a weaker form and without explicit bounds in Graph Minors XIII by Robertson and Seymour [79]. The best known bounds are due to Chuzhoy [11]. The version below is due to Gorsky, Seweryn, and Wiederrecht [38].
Proposition 2.4 (Flat Wall Theorem [79, 38]).
Let be integers with let and let be a graph with an -wall Then either
there exists a model of in which is controlled by or
there exist a set with and an -subwall of which is disjoint from and flat in
Furthermore, there exists a -time algorithm which finds either the model of or the set the subwall and a -rendition witnessing that is flat in
A slightly weaker but relevant version of grounded walls is the notion of centrality. Let be a surface and let be a -rendition of a graph containing an -wall We say that is -central if there is no cell such that contains the -majority side of a separation from Similarly, let let be a surface and be a -rendition of Then we say that is -central for if no cell of contains the majority side of a separation from
With this, we are ready to state the so called Local Structure Theorem. This theorem was originally proven by Robertson and Seymour [72], the version we present here is the recent result of Gorsky, Seweryn, and Wiederrecht [38] that also allows for polynomial bounds.
Proposition 2.5 (Local Structure Theorem [72, 38]).
There exist functions and such that for all integers, and every graph on vertices, every graph and every -wall one of the following holds.
i)
has an -minor controlled by or
ii)
there exists a set of size at most and a subwall such that has a -rendition of breadth at most and depth at most such that is a surface where does not embed, is flat in and is -central.
Moreover, it holds that There also exists an algorithm that, given a graph a graph and a wall as above as input finds one of these outcomes in time
2.4 Well-Quasi Ordering
The main goal of this subsection is to prove that for fixed the colorful minor relation on colorful graphs on at most colors is a well-quasi-order. This result follows directly from a powerful general theorem of Robertson and Seymour [74].
In this subsection, and this subsection only, we permit our graphs to have loops.
A directed graph is an oriented graph if there exists a graph such that and can be obtained from by replacing each undirected edge by one of the two possible directed edges or
Proposition 2.6 (Robertson and Seymour [74]).
Let be a well-quasi-order and be a sequence of oriented graphs. Moreover, for each let be some function. Then there exist and a map with domain satisfying:
i)
for each is a connected subgraph of and there exists with ; and for all pairs of distinct
ii)
for each is an edge of a loop if and only if is a loop, with ; and for all distinct pairs and
iii)
for each with head and tail has head in and tail in and is not an edge of or (the last is trivial unless ).
This is all we need to deduce theorem˜1.1.
Proof of Theorem˜1.1..
We consider the partial order Since is a fixed integer, is finite and thus, is a well-quasi-order.
Suppose there exists an infinite but countable anti-chain of colorful graphs on at most colors under the colorful minor relation. For each we define for all and for all Moreover, for each we may assign an arbitrary orientation to the edges of to obtain an oriented graph.
Then there must exist such that and satisfy the outcome of proposition˜2.6 witnessed by the map Now observe that also witnesses that is a minor of and, in particular, for all Hence, is a colorful minor of which contradicts the assumption that is an anti-chain. ∎
Colorful obstructions.
Let be a non-negative integer and let be a class of -colorful graphs that is closed under taking colorful minors. We define the (colorful minor) obstruction set of denoted by as the set of all colorful minor minimal -colorful graphs that are not members of Notice that the obstruction set of is an anti-chain for the colorful minor relation, therefore, from theorem˜1.1 we have the following.
Corollary 2.7.
Let be a non-negative integer and let be a class of -colorful graphs that is closed under taking colorful minors. Then is a finite set.
2.5 Colorful minor checking
We now proceed with the proof of theorem˜1.6. Having established that the colorful minor relation is indeed a well-quasi-order, we next show that testing whether some -colorful graph is a colorful minor of a -colorful graph can be done efficiently. Towards proving theorem˜1.6 we require the fast minor testing algorithm of Korhonen, Pilipczuk, and Stamoulis [56]. Indeed, we need the rooted variant of their algorithm.
A tuple where for all is called a rooted graph and the are called the roots of . Notice that one may interpret a rooted graph with roots as an -colorful graph where for each . Indeed, to each rooted graph we associate a corresponding colorful graph such that for all . We say that a rooted graph is a rooted minor of a rooted graph if is a colorful minor of .
Proposition 2.8 (Korhonen, Pilipczuk, Stamoulis [56]).
There exists a function and an algorithm that takes as input two rooted graph and and decides if is a rooted minor of in time .
The proof of theorem˜1.6 is a simple reduction of the colorful minor checking problem to the rooted minor checking problem.
Let be a -colorful graph. We denote by the rooted graph obtained from by introducing for every and a new vertex adjacent precisely to .
Moreover, let be another -colorful graph. We denote by the rooted graph obtained from by introducing, for every and a new vertex adjacent to all vertices of .
Notice that for a fixed -colorful graph , the number of roots in and are the same. As a convention, we will assume the vertices of to come with some linear ordering such that in the sequence of roots of , the vertex has precisely the same position as the vertex in the sequence of roots of . Under this convention, we now obtain the following lemma.
Lemma 2.9.
Let and be -colorful graphs. Then is a colorful minor of if and only if is a rooted minor of .
Proof.
Assume first that there exists a colorful minor model of in . In this case, for each and there exists a vertex with . Hence, in , is adjacent to . It follows that is a minor model witnessing that is indeed a rooted minor of .
Now assume that there exists a minor model of in . In this case, for each , must first of all be disjoint from all roots of as each must be mapped by the minor model to the root of . Moreover, for each , there is a vertex of adjacent to . It follows from the definition of that . Hence, must be a colorful minor model of in . ∎
With lemma˜2.9 it now suffices to observe that both and can be computed from and in time linear in . Moreover, the number of roots of is at most and this is also an upper bound on the number of vertices has in excess of itself. Finally, the number of edges in is at most . So we may compute and run the algorithm from proposition˜2.8 to check for as a rooted minor without any additional impact on its running time in terms of and .
3 Excluding a rainbow clique
In this section we discuss the structure of colorful graphs excluding a fixed rainbow clique as a colorful minor. For this, first we establish a general tool that allows us to connect several sets of vertices to a fixed target set with many disjoint paths, or to find a small separator that separates at least one of the from From here, we make use of several useful results of Robertson and Seymour from [79] to establish the following key theorem of this section.
Theorem 3.1.
Let and be positive integers with Let be a -colorful graph such that contains a minor model of Then one of the following is true.
i)
There exists a colorful minor model of a rainbow in such that is a truncation of or
ii)
There exists a set of size at most such that the -big component of is restricted.
Furthermore, there exists an algorithm that takes as input and as above and finds one of the two outcomes in time
Establishing theorem˜3.1 is the key towards our first structure theorem for colorful graphs excluding a rainbow clique minor, i.e. theorem˜1.2. Robertson and Seymour in [79], prove a theorem akin to theorem˜3.1 that however works in the setting of -colorful graphs with a bounded number of colored vertices. In the following section, we adapt this result to work for colors and an arbitrary number of colored vertices to suit our needs.
3.1 Finding a rainbow clique
Within our proofs, in this section and beyond, we will repeatedly make use of a variant of proposition˜2.1 that allows us to link a family of a bounded number of vertex sets equally into a single set or separate at least one of the from by deleting a small set of vertices.
Lemma 3.2.
For all positive integers every graph and every collection of (not necessarily disjoint) vertex sets one of the following holds:
i)
there exists a linkage of order in such that for every there exists an --linkage of order and or
ii)
there exists a non-empty set and a set of order less than such that is a --separator in
Moreover, there exists an algorithm that, given and as above as input finds one of the two outcomes in time
Proof.
Let and be given. We construct an auxiliary graph as follows. For every we introduce a set of new vertices and for each we make adjacent to all vertices of We then apply proposition˜2.1 to the graph with sets and to either find a --linkage of order or a --separator of order at most in
Let us first assume the case where proposition˜2.1 returns the linkage Then every vertex of must be an endpoint of some path in For each let the collection of paths obtained from a path with one endpoint in by deleting this endpoints from It follows that and is an --linkage. Hence, this case produces a linkage as desired.
So now assume proposition˜2.1 returns a --separator of order at most in Let be the set of all such that there is no --path in All we have to do is to show that is non-empty. Towards this goal let be the collection of all such that is non-empty. Since and we know that Now observe that there cannot be a such that there is an --path in To see this assume that such a path, let us call it exists for some Then this path must also exist in However, since there is some and has one endpoint in the path is a --path in This contradicts the assumption that is a --separator in Thus, the set is the set required by the second outcome of our assertion and the proof is complete. ∎
We are going to need the following theorem of Robertson and Seymour [79].
Proposition 3.3 (Robertson and Seymour [79]).
Let and be positive integers with Further, let be a graph, of size and be pairwise vertex-disjoint subgraphs of such that
i)
for every either is connected, or every component of has a vertex in
ii)
for all either both and contain a vertex of or there exists an edge with and and
iii)
there is no such that there is a separation of of order less than with and
Then there exists together with connected and pairwise vertex-disjoint subgraphs of such that
for each and for all and
for all if there does not exist an edge with and then i.e. both and have a non-empty intersection with
Notice that proposition˜3.3 can be realized in polynomial time, allowing for an algorithmic version. From here on, all of our proofs are constructive and thus, all of them admit polynomial-time algorithms that yield the respective outcomes.
The following does not find a direct application in this paper, but it makes for a nice observation on its own.
Observation 3.4.
Let and be positive integers. Let be a -colorful graph such that there exist pairwise vertex-disjoint connected subgraphs such that for all there exists with and Then there exist and such that
and
for all it holds that
Proof.
The lemma follows immediately from an iterative application of the pigeon hole principle as follows. Let For each assuming inductively that with has been found, let and Then there exists such that if this is true for both set Let Then, by construction, the set has the desired properties. ∎
Lemma 3.5.
Let and be positive integers. Let be a -colorful graph such that there exist connected and vertex-disjoint subgraphs of such that for every there exists an edge such that and Then one of the following is true.
i)
there exists a separation of of order less than a non-empty set such that and there exists such that or
ii)
for there exist pairwise vertex-disjoint subgraphs such that
for every and it holds that
for all either is connected or there exists such that and every component of contains a vertex with and
for every there either exists an edge such that and or
Proof.
Let us begin by creating an auxiliary graph as follows. For every introduce a set of new vertices such that each of them is adjacent to every vertex with Let Then
Now assume that there exists a separation of of order at most such that and there is some such that Let and let We claim that there exists some such that and is a separation of order at most of First notice that, if there is an edge with and then this edge also exists in Hence, is indeed a separation of of the desired order. Next observe that, since and there must exist some such that It follows from the definition of separations that there cannot be a vertex such that as otherwise would be adjacent to a vertex in Thus, in this case the first outcome of our assertion holds.
From here on we may assume that there does not exist a separation in such that and there is some such that In this situation we may apply proposition˜3.3 to and the graphs It follows that there exists some together with connected and pairwise vertex-disjoint subgraphs of such that
for each and for all and
for all if there does not exist an edge with and then
By possibly reordering we may assume that for every and every it holds that It then follows from the definition of that for each For every let Now, for every and it holds that either is connected, or every component of contains a vertex that is adjacent to a vertex of in It follows that this vertex carries color Finally, the last point of outcome two of our assertion follows directly from the definition of the and our proof is complete. ∎
Proof of Theorem˜3.1..
Let We may now apply lemma˜3.5.
Suppose lemma˜3.5 returns a separation of of order at most together with a non-empty set such that and there is some for which It follows that is restricted. Moreover, the unique -big component of must be fully contained in Hence, with we have found the desired set.
Hence, we may assume that lemma˜3.5 returns pairwise vertex-disjoint subgraphs where and
for every and it holds that
for all either is connected or there exists such that and every component of contains a vertex with and
for every there either exists an edge such that and or
Now, for each let us define a connected graph as follows. For each there exists a component of such that and there exists an edge with and We set It is now easy to see that is indeed a colorful minor model of a rainbow ∎
3.2 Colorful graphs without large rainbow clique minors
We now have everything in place to prove theorem˜1.2. In fact, we will prove a slightly stronger statement which will then imply theorem˜1.2.
Theorem 3.6.
There exist functions such that for all integers and every -colorful graph one of the following holds:
i)
contains a rainbow as a colorful minor, or
ii)
has a tree-decomposition of adhesion at most and a (possibly empty) subset of the leaves of such that for every one of the following holds:
is a leaf with parent and is restricted, or
if is the colorful torso of at then there exists a set of size at most and a surface of Euler-genus at most such that has a -rendition of breadth at most and width at most
Moreover, it holds that and
There also exists an algorithm that, given and as input, finds either a colorful minor model of a rainbow in or a tree-decomposition as above in time
Our proof of theorem˜3.6 proves a slightly stronger variant that allows for a straightforward induction.
Let be a graph. A tuple is a rooted tree-decomposition for if is a tree-decomposition for and We also define notions like width and adhesion for a rooted tree-decomposition by forwarding them to the tree-decomposition
Theorem 3.7.
There exist functions such that for every every every -colorful graph and every vertex set with one of the following holds:
i)
contains a rainbow as a colorful minor, or
ii)
has a rooted tree-decomposition of adhesion at most and a (possibly empty) subset of the leaves of such that for every one the following holds:
is a leaf with parent and is restricted, or
if is the colorful torso of at then there exists a set of size at most and a surface of Euler-genus at most such that has a -rendition of breadth at most and width at most In addition, if then
Moreover, it holds that
There also exists an algorithm that, given and as input, finds either a colorful minor model of a rainbow in or a tree-decomposition as above in time
Proof.
We begin by declaring the functions involved. Let be the constant from proposition˜2.3.
Notice that the required bounds on all four functions follow directly from proposition˜2.5 and the above definitions.
We now proceed by induction on and start by discussing two fundamental cases. In the following we assume that does not contain a rainbow as a colorful minor.
Fundamental case 1: In case we may select to consist solely of the vertex and set which satisfies the second outcome of our theorem. Thus, we may assume
Fundamental case 2: Similarly, if we may select any vertex and set thereby achieving and so we are done by induction. Thus, we may assume
We now call proposition˜2.2 for and which has two possible outcomes. Either, is -linked, or there is a -balanced separator of size at most for We treat these two cases separately.
Case 1: There exists a -balanced separator of size at most for
In this case, for each component of we consider the graph together with the set Notice that, since is a -balanced separator for of size at most we have that
Hence, and satisfy the properties of one of the two fundamental cases. In either case, we obtain a rooted tree-decomposition together with a set of leaves for with From here we define and as follows. The tree is the tree obtained from the disjoint union of all by introducing a new vertex adjacent to all The bags of are set to be in case for some component of and otherwise we have and set Moreover, we set to be the union of all Notice that is indeed a rooted tree-decomposition as required and, in particular, and so satisfies the second bullet point of the second outcome of our claim.
Case 2: The set is -linked.
Since we may now apply proposition˜2.3 to and obtain a -wall such that This allows us to apply proposition˜2.5 to asking for a clique minor of size There are two possible outcomes:
i)
contains a -minor controlled by or
ii)
there exists a set of size at most and a subwall such that has a -rendition of breadth at most and depth at most such that is a surface where does not embed, is flat in and is -central.
We further distinguish two subcases based on these outcomes.
Case 2.1: The graph contains a model of controlled by
In this case, we apply theorem˜3.1 to in Since, by our assumption, does not contain a rainbow as a colorful minor, the second outcome of theorem˜3.1 must hold. That is, we obtain a set of size at most such that the -big component of is restricted.
Now let be any other component of apart from Let Since is the -big component of and as is controlled by we obtain that
Hence, for each component of other than we have that together with falls into one of the two fundamental cases. Therefore, for each such we obtain a rooted tree-decomposition and a set of leaves of with that satisfies the second outcome of our claim.
We define the desired rooted tree-decomposition with the set of leaves as follows. Let be the tree obtained from the disjoint union of the for all components of by introducing a new isolated vertex and then a new vertex adjacent to and all We set and for all such that there is a component of with Finally, we set to be the union of all together with the set As before, it is straightforward that together with is indeed the required rooted tree-decomposition as demanded by the second outcome of our claim. This settles the first subcase.
Case 2.2: There exists a set of size at most and a subwall such that has a -rendition of breadth at most and depth at most such that is a surface where does not embed, is flat in and is -central.
Since is -central, no cell of contains the majority side of any separation in Since it follows that no cell of contains the big side of a separation from Therefore, we have for all non-vortex cells Moreover, for every vortex and every bag from the linear decomposition of adhesion at most – which is guaranteed to exist due to our assumptions on – we have that also
For every non-vortex cell let and Similarly, for every vortex cell and every where is the number of bags in the linear decomposition of let and be the union of the – at most two – adhesion sets of and the vertex of in It follows that
for all non-vortex cells all vortex cells and all
Now for each and each as before, we may observe that, together with their respective and they fall into one of the two fundamental cases. From here, we obtain for each of them, a rooted tree-decomposition and a set ( and for the vortices and respectively) as desired. Finally, all that is left to do is to combine them into the rooted tree-decomposition with set as follows.
Let be the tree obtained from the disjoint union of all and by adding a new vertex adjacent to the and We set where denotes the union of all the adhesion sets of the linear decompositions of all of the vortices From our inequalities above it follows that
as anticipated. For all such that there is either a non-vortex cell with or a vortex cell and some such that we set and respectively. Finally, the set is the union of all sets and With this, our proof is complete. ∎
4 Excluding a rainbow grid
We now move on from rainbow cliques to rainbow grids: In this section, we explain how theorem˜1.4 can be deduced from a powerful variant of proposition˜2.5 by Paul, Protopapas, Thilikos and Wiederrecht [69, 68].
4.1 Colors in a surface
The main purpose of this subsection is to import a set of tools from the work of Paul, Protopapas, Thilikos and Wiederrecht in [69] (see also [68]). These tools were developed to homogenize -renditions with respect to the rooted minors one can find within their cells. In [69] rooted minors are represented using colors, a feature that comes in handy for our purposes.
Colorful Boundaried Indices.
Consider a fixed non-negative integer A boundaried -colorful graph is a quadruple where is a -colorful graph, and is a bijection. In other words is an alternative compact notation for a rooted -colorful graph where all roots are pairwise distinct, i.e., for the rooted -colorful graph
Two boundaried -colorful graphs and are compatible if is an isomorphism from to (where ), and for each Moreover, they are isomorphic if there exists an isomorphism between and such that and for every We denote by the set of all (pairwise non-isomorphic) boundaried -colorful graphs whose boundary contains at least and at most vertices. A -colorful boundaried index is a function such that, if for two -colorful boundaried graphs and then and are compatible. Moreover, we demand that if then We define the capacity of as We moreover demand that for some positive integer
Index representation.
For the purposes of this section, and this section only, we assume that every -colorful graph comes equipped with a universal ranking, i.e. an injection This is useful as it allows us to uniquely define boundaried graphs on its subgraphs. For instance, if is a subgraph of and then we may define the boundaried graph where for every if and only if is the -th smallest integer in This allows us to utilize the simpler notation instead of
Next, we would like to express what it means for a collection of indices to be bidimensionally represented within a -rendition. This is exactly what will allow us to obtain the rainbow grid later on.
Let be a -colorful boundary index, be a -colorful graph, and be a -rendition of Given we say that a minor model of a -grid in represents in if
is disjoint from for all vortices and
for every vertex for every if there exists a non-vortex cell with then there is a cell with and Here denotes the value
An indexed local structure theorem.
We are now ready to state a key result from [69]. A preliminary version was already used in [68], but [69] contains the following, much more streamlined version.
Proposition 4.1.
There exists functions and such that, for every non-negative integer for every -colorful boundary index with every every every graph on vertices, every -colorful graph and every -wall one of the following holds:
i)
contains as a minor controlled by or
ii)
there is a set a -rendition of and a -wall such that
is a surface where does not embed,
has breadth at most and depth at most
is grounded in and and
there is a minor model of a -grid in that represents in
Moreover,
Also, given an algorithm for there exists an algorithm that, given a graph a graph and a wall as above as input, finds one of these outcomes in time
We should stress that the original formulation of proposition˜4.1 is not stated in terms of -colorful graphs and -colorful boundary indices. However, a variant like proposition˜4.1 is straightforward to obtain as in the original proof the boundary index is only used to assign a coloring to the non-vortex cells of a -rendition of which in this case would depend on Besides that, would not play any other role in its proof and can safely be disregarded thereafter.
4.2 Locally excluding a rainbow grid
Proposition˜4.1 allows us to immediately deduce a strong local structure theorem for -colorful graphs excluding rainbow grids.
For our purposes, we make use of a highly simplified case of a colorful boundaried index. We wish to express -colorful graphs as -colorful boundaried graphs in a way that preserves the information of colors the vertices on the boundary have access to. We have to be slightly careful so as to not map two non-equivalent -colorful boundaried graphs to the same index even if their boundaries have access to the same set of colors.
To achieve this, let be a partition of to maximal sets of equivalent -colorful boundaried graphs such that for every if and Notice that since we consider boundaries of size at most is bounded by a constant depending only on Consider a bijection such that for every and every
Now, for every non-negative integer we define a -colorful boundaried index so that for every -colorful boundaried graph for some where is the component of that contains In other words, expresses the set of all colors that can be represented by one of the boundary vertices.
Lemma 4.2.
There exist functions such that, for all non-negative integers and every every -colorful graph and every -wall one of the following holds:
i)
contains the rainbow -grid as a colorful minor,
ii)
there exists a set of size at most such that the -big component of is restricted, or
iii)
there exists a set a -rendition of and a -wall such that
is a surface of Euler-genus in
has breadth at most and depth at most
is grounded in and and
if there exists such that for all non-vortex cells
Moreover,
Also, there exists an algorithm that, given a -colorful graph and a wall as above as input, finds one of these outcomes in time
Proof.
First, we set up our functions as follows:
for a non-negative integer which depends only on and such that
We proceed by applying proposition˜4.1 with and Notice that this means that This yields one of two outcomes as follows.
contains a -minor controlled by
Let be the minor model of and notice that is a truncation of We now apply theorem˜3.1. If this results in a colorful minor model of a rainbow we have found every -colorful rainbow graph on vertices as a colorful minor in and thus, in particular, we have found the rainbow -grid. Hence, we may assume that we enter the second outcome of theorem˜3.1 which yields a set of size at most
such that the -big component of is restricted. Since this means we have reached the second outcome of our assertion and thus we are done.
does not contain a -minor controlled by
In this case we now find a set of size at most a -wall whose tangle is a truncation of the tangle of a surface where does not embed, and a -rendition of breadth at most and depth at most such that is grounded in and bidimensionally -represents
Notice that the Euler-genus of is in as desired.
It follows that there exists a minor model of the -grid in which is disjoint from the interiors of the vortices of such that for each and every for which there is a non-vortex with there exists a non-vortex cell with such that
Notice that the boundary of any non-vortex cell of acts as a separator of size at most in Hence, for any and each non-vortex cell with it follows that each component of that contains a vertex from must be vertex-disjoint from all Hence, for each such we may assume and, moreover, since is connected, this means that
Suppose for every there exists and such that and for some Let us denote by the set of all vortex cells of Then we have that
Since represents in this implies that for all which means that is indeed a colorful minor model of the rainbow -grid.
Let
By our discussion it follows that, if and contains the rainbow -grid as a colorful minor which is the first outcome of our assertion. Hence, we may assume Moreover, for every we may assume that every component of contains a vertex from as we may remove any component of that does not contain a vertex of and add it to for some without altering the essential properties of In the case where we may artificially introduce a new cell that acts as the storage for all such components.
Therefore, if we have that and for all non-vortex cells as desired. This completes the proof. ∎
4.3 The proof of Theorem˜1.4
We are now ready for the proof of Theorem˜1.4. The proof will largely follow a similar strategy as the one for theorem˜1.2. Indeed, we will prove a stronger version similar to theorem˜3.6 and its strengthened form theorem˜3.7. Theorem˜1.4 will then follow as a corollary. This slightly stronger version reads as follows:
Theorem 4.3.
There exist functions such that for all and all -colorful graphs one of the following holds:
i)
contains the rainbow -grid as a colorful minor, or
ii)
has a tree-decomposition of adhesion at most such that for all one the following holds:
is a leaf with unique neighbor such that is restricted, or
if is the colorful torso of at then there exists a set of size at most a surface of Euler-genus at most and a non-empty set such that has a -rendition of breadth at most and width at most such that for all non-vortex cells and if and only if for some vortex
Moreover,
Also, there exists an algorithm that, given and a -colorful graph as input, finds one of these outcomes in time
As announced, we strengthen theorem˜4.3 further to allow for a smooth induction.
Theorem 4.4.
There exist functions such that for all all -colorful graphs and every set with one of the following holds:
i)
contains the rainbow -grid as a colorful minor, or
ii)
has a rooted tree-decomposition of adhesion at most and a (possibly empty) subset of leaves of such that and for all one the following holds:
is a leaf with a unique neighbor such that is restricted, or
if is the colorful torso of at then there exists a set of size at most a surface of Euler-genus at most and a non-empty set such that
a)
has a -rendition of breadth at most and width at most such that for all non-vortex cells and if and only if for some vortex and
b)
if is a child of with for some non-vortex cell then
a)
Moreover,
Also, there exists an algorithm that, given and a -colorful graph as input, finds one of these outcomes in time
Proof.
First, we define the functions involved as follows where is the constant from proposition˜2.3.
Moreover, in the following let
We now proceed by induction on and start by discussing two fundamental cases. Exactly as in our proof of theorem˜3.7, we first discuss the two fundamental cases where either has only few vertices, or These two cases are fully identical to the ones in theorem˜3.7 and we therefore omit them here. Then two further cases follow:
Case 1, where has a small balanced separator, and Case 2, where does not have a small balanced separator. As before, for Case 1 there is no difference to the proof of theorem˜3.7 and thus, we may omit it here.
This leaves us with the last case.
Case 2: The set is -linked.
Since we may now apply proposition˜2.3 to and obtain a -wall such that This further allows for the application of lemma˜4.2 to the wall in In the following we distinguish three cases based on the three possible outcomes of lemma˜4.2.
Case 2.1: contains the rainbow -grid as a colorful minor.
This is one of the possible outcomes of our assertion and thus, there is nothing more to show.
Case 2.2: There exists a set of size at most such that the -big component of is restricted.
Let be the -big component of and let denote all other components of Observe that, since we have that And thus, for all
We now describe how to create the rooted tree-decomposition as required by the assertion.
First, for each select a vertex and let Notice that, by definition and therefore we have that and, moreover, for all Thus, we may apply the induction hypothesis to each thereby obtaining a rooted tree-decomposition with leaf sets as described in the assertion for each In particular this means
Let now be the tree obtained from the disjoint union of the by introducing a vertex and a vertex where is adjacent to and all Moreover, we set We set and as well as for each and every One can now easily observe that is the desired rooted tree-decomposition of
Case 2.3: There exists a set a -rendition of and a -wall such that
is a surface of Euler-genus in
has breadth at most and depth at most
is grounded in and
there exists such that for all non-vortex cells
Let us begin with a number of minor observations. First of all, is grounded in This means that for every non-vortex cell no entire row or column of can be contained in Since has size at most and separates from the rest of in and it follows that where the additional vertices may only be contained in
Similarly, let be any vortex with linear decomposition of adhesion at most It follows that there exists a set of size at most that separates from the rest of Moreover, is disjoint from Indeed, one can observe that no column or row of may be fully contained in Hence, we deduce that
Now, for every non-vortex cell let select if it exists, and let Notice that and or Moreover, from the fact that we may deduce that Hence, each of the is restricted.
Similarly, for every vortex and let select if it exists and let
As before, one may observe that and or
These observations give us the right to call the induction hypothesis for every pair of a vortex and on and This yields, in the absence of a rainbow -grid as a colorful minor, rooted tree-decompositions as described by the assertion for each such pair
We are now ready to construct the rooted tree-decomposition as follows. Let be the tree obtained from the disjoint union of the for all and all vortices by introducing, for each non-vortex cell a new vertex then introducing a new vertex and finally joining to all and all We set
For each non-vortex cell we set and for every vortex and each we set for all It is now easy to check that is indeed a tree-decomposition and from our discussion above it follows that it has exactly the properties required by our assertion. ∎
5 Excluding segregated grids
In this section, we are concerned with the proof of theorem˜1.3, which is an analogue of the Grid Theorem of Robertson and Seymour for torso treewidth. Similar to the organization of section˜4, we begin with some preliminary results and a “local” variant of the structure theorem before proceeding to the actual proof of theorem˜1.3.
5.1 A homogeneous flat wall
In this section we prove an auxiliary result on flat walls which has been proven in several variations before. We include our own proof specifically for the setting of colorful graphs for the sake of completeness.
Let be a -colorful graph with a flat -wall where and a -rendition witnessing that is flat in and where is the sphere. Let be a brick of Then let be the minimal disk such that contains but not the perimeter of Notice that by the minimality of it follows that is the only brick of contained in
We say that is homogeneous if for all bricks and of we have
Lemma 5.1.
There exists a function such that for all integers and and every -colorful graph and every -wall in one of the following holds:
i)
there exists a model of in which is controlled by or
ii)
there exists a set with and an -subwall of which is disjoint from and flat and homogeneous in
Moreover, and there exists a -time algorithm which finds either the model of or the set the subwall and a -rendition witnessing that is flat in
Proof.
We begin by defining an auxiliary recursive function as follows.
From this recursion it follows that We now set
This allows us to call upon proposition˜2.4 to find either a model of controlled by in which case we would be done, or a set of size less than and an -subwall that is flat in
From here on, we will be further refining For now let be the – possibly empty – collection of all colors such that for all bricks of and let the – also possibly empty – collection of all colors such that there does not exist a single brick of with
Our goal is to find sets and together with a subwall of of order at least such that is homogeneous, for all bricks of and This is done by induction over the number of colors and in each step we add precisely one new color to either the set or the set until all colors have been distributed. Clearly, such a process must always stop after at most steps. Moreover, the base case is trivially true.
Suppose for we have already constructed the sets and as well as an -wall such that and for all bricks of and Then fix to be as small as possible. In case no such exists we are already done, so we may assume that there is a possible choice.
Next, let us denote the vertical paths of by and the horizontal paths of by Next, for each let Let be the -wall consisting of the vertical paths and horizontal paths for all Observe that for every brick of the graph contains an -subwall of whose perimeter is precisely the cycle
Now we are in one of two cases: Either for all bricks of or there exists a brick of such that
In the first case we let be the -subwall of that avoids its perimeter, and set and In this case we are done by induction.
In the second case, we let be the -subwall of that avoids and we set and Hence, also here we are done by induction and our proof is complete. ∎
5.2 Locally excluding segregated grids
We are now ready to provide a full local structure theorem for -colorful graphs that exclude all -segregated grids as colorful minors.
Theorem 5.2.
There exist functions such that for all integers and every -colorful graph and every -linked set one of the following holds:
i)
contains a -segregated grid as a colorful minor, or
ii)
there exists a set with such that the -big component of is restricted.
Moreover, and there exists an algorithm that, given and as above as input finds one of these outcomes in time
Proof.
As it is custom, we begin by setting up our numbers. Let be the constant from proposition˜2.3 and hfw be the function from lemma˜5.1.
As a first step, we apply proposition˜2.3 to and obtain a -wall Next, we apply lemma˜5.1 to This either results in a -minor model with or in a set of size at most together with a homogeneous flat wall in of order
We treat these two cases separately.
Case 1: We find a -minor model with
We now apply theorem˜3.1. This either yields a rainbow as a colorful minor in or a set of size at most such that the -big component of is restricted. With being a truncation of the second outcome matches the second outcome of our assertion. Hence we may assume to find a rainbow as a colorful minor. This, however, means that contains every -colorful graph on vertices as a colorful minor and this includes all -segregated grids. Hence, we are done with this case.
Case 2: We obtain the set of size at most together with a homogeneous flat wall in of order
We now layer our wall as follows. Notice that for any -wall with perimeter contains a -wall that is a subwall of If for some positive we may repeat this process.
Let now be the first cycles obtained from by iteratively removing perimeters as described in the process above and let be the remaining -wall. We assume the cycles as above are numbered in order of their removal, that is, is the perimeter of If we number the horizontal paths of starting from the top one as then the horizontal paths of are subpaths of Let us select to be the set of the first vertex on from each with when traversing along starting on its left-most endpoint.
Now let for some brick of Since is homogeneous, we know that also is homogeneous, so the choice of here is arbitrary. Let and for each let us denote by the set We now apply lemma˜3.2 to the sets and the set asking for many paths for each member of This yields one of two outcomes. Either we find a linkage of order such that contains a --linkage of order for each and the paths in are all internalle vertex-disjoint from or there is a set of size less than such that there exists some and there is no --path in
We discuss these two outcomes independently.
Case 2.1: There exists a linkage of order such that contains a --linkage of order for each and
In this case we consider the left most vertical path of which contains all vertices of We then partition into disjoint paths as follows. Let be the shortest subpath of containing the top-most endpoint of such that there exists and there are paths in with an endpoint in Notice that for all at most paths from have an endpoint on
Now suppose, for some we have already found disjoint and consecutive subpaths of such that there is an injection and for every paths from have an endpoint on Moreover, we may assume that for each with for all at least paths of have an endpoint on
Next, let be the shortest subpath of containing a neighbor of such that there exists for all and paths from have their endpoints on Clearly, by our choices of the so far, such a path and color exist. Moreover, if we set we now have consecutive subpaths of such that for each paths from have their endpoint in now is an injection from to and for all remaining colors in there are at least endpoints of paths from on
Hence, we find a bijection from to and linkages of order for each such that for every all paths in have their endpoint on Let now be the -subwall of obtained by removing, similar to before, all but layers from Notice that these layers give rise to a family of cycles where
Moreover, there are pairwise vertex-disjoint paths – subpaths of horizontal paths of – that connect the left most vertical path of to the left most vertical path of and which have all their internal vertices in Let now be the set of all endpoints of the paths in Moreover, let be the set of endpoints of the top most horizontal paths of and the left most vertical path of
We claim that there exists an --linkage of size such that all internal vertices of the paths of belong to To see this, notice that in the absence of such a linkage proposition˜2.1 would imply the existence of a separation of order at most between and However, each vertex of belongs to a horizontal path of that can be extended in at least different ways, intersecting only in the horizontal path, to a path that intersects all cycles at least of Hence, such a separator must leave at least one cycle of one way to get from some vertex of to and one way to get from to intact. Hence, deleting any set of size at most still leaves at least one --path intact and therefore, and cannot be separated by fewer than vertices in
Let be the perimeter of and be the -subwall of that avoids It is now easy to see that together with the paths from the paths from the horizontal paths of and the bricks of that intersect we may construct a -segregated grid as a colorful minor of See figure˜4 for an illustration.
Case 2.2: There exists a set of size less than such that there exists some and there is no --path in
In this case we claim that the -big component of is restricted. To see this, simply notice that we removed cycles from to create Moreover, there exist pairwise vertex-disjoint horizontal paths of that contain a horizontal path of and therefore a vertex of Since in there must still be a path from the perimeter of to Indeed, take any pair of one horizontal and one vertical path of that is disjoint from then there exists a path from this pair to in the -big component of Moreover, there does exist such a path which is internally vertex-disjoint from Thus, if the -big component of would contain a vertex with then there would still be a --path in which is impossible. Hence, in this case the second outcome of our assertion holds. ∎
5.3 A Grid Theorem for torso treewidth
We are now ready for the proof of theorem˜1.3. Similar to our proofs of Theorems˜3.6 and 1.4 we do this by proving a slightly different and stronger statement which, in turn, is implied by a further strengthening that lends itself better to induction.
Theorem 5.3.
There exists a function such that for all positive integers and and every -colorful graph one of the following holds:
i)
contains a -segregated grid as a colorful minor and the torso treewidth of is at least or
ii)
there exists a tree-decomposition for such that there is a set of leaves of and
has adhesion at most
for all and
for each with neighbor is restricted.
Moreover, and there exists an algorithm that, given and as input, finds either a colorful minor model of a -segregated grid in or the tree-decomposition as above in time
Notice that the tree-decomposition from the second outcome of theorem˜5.3 is a witness for the torso treewidth of being at most Hence, theorem˜5.3 does indeed act as an analogue of the Grid Theorem of Robertson and Seymour for torso treewidth.
We first prove that -segregated grids indeed act as lower bounds on the torso treewidth of a graph.
Lemma 5.4.
For all positive integers and and every -colorful graph if contains a -segregated grid as a colorful minor, then the torso treewidth of is at least
Proof.
Suppose, towards a contradiction, that contains a minor model of some -segregated grid and, at the same time, that there exists a set such that where denotes the torso of in
Let be a tree-decomposition of width at most for and let For each let be the component of containing and denote by the set It follows directly from the definition of tree-decompositions that is a separation of order at most in
Recall that for every clique of a graph and every tree-decomposition of there must exist such that Hence, for each component of there exists such that This observation allows us to make a fundamental assumption on the structure of From now on we will assume that for every component of there exists a leaf such that
For each let denote the union of and the vertex sets of all components of such that Then is a separation of order at most in
Since is a minor model of the -grid in particular it follows that is a tangle of order in Hence, we may assign to each edge with the orientation and thereby obtain an orientation of
Suppose first that for every leaf of with unique neighbor we have In this case there must exist a unique sink that is not a leaf in Notice that this means, in particular, that for all pairs of components of if and belong to different components of then and are not connected in This is, because by the tangle axioms it follows that no vertex of can have two outgoing edges. Let be the neighbors of Notice that for every it holds that Hence, if we have that which contradicts being a tangle. So we may assume that
We prove by induction on that for any subset of order at most we have For this is trivially true by choice of and the definition of So assume the claim holds for all and let be any set of size In case we have that we know, by induction that for any we have
Moreover, it holds that
Hence, we obtain that
which is a contradiction to being a tangle.
Consequently, we may now assume that there exists a leaf which is the unique sink of In case there does not exist a component of such that we may apply a similar argument as above and obtain a contradiction to being a tangle. Hence, there exists a component of such that Let be the unique neighbor of Then is the -big component of Thus, there exists some vertical path of a -segregated grid such that there is a path in which meets each with
Indeed, it follows that for each there exists a set of vertices with such that there is a --linkage of size in Hence, for every color there exists a --linkage of size in However, by our assumption we have that is restricted and Since this is clearly absurd we have found our final contradiction and our proof is complete. ∎
All that is left to do is to prove the following strengthening of the upper bound part of theorem˜5.3.
Theorem 5.5.
There exists a function such that for all positive integers and every -colorful graph and every set of size at most one of the following holds:
i)
contains a -segregated grid as a colorful minor, or
ii)
there exists a rooted tree-decomposition for such that there is a set of leaves of and
has adhesion at most
for all and
for each there exists a neighbor of such that is restricted.
Moreover, and there exists an algorithm that, given and as input, finds one of these outcomes in time
Proof.
We quickly go through the part of the proof that is analogous to previous variants of a “local-to-global”-type argument.
First, the function is defined as follows.
We now proceed by induction on and start by discussing two fundamental cases. Exactly as in our proof of theorem˜3.7, we first resolve the two fundamental cases where either has only few vertices, or in precisely the same way as before. Then two further cases follow:
Case 1, where has a small balanced separator, and Case 2, where does not have a small balanced separator. As before, for Case 1 there is no difference to the proof of theorem˜3.7 and thus, we immediately move on to Case 2.
This leaves is with the last case.
Case 2: The set is -linked.
In contrast to previous iterations of this proof, here we do not have to distinguish more subcases. We immediately may apply theorem˜5.2 to and which either yields a -segregated grid as a colorful minor in – and there would be nothing more to prove – or it provides us with a set with such that the -big component of is segregated.
In this second case we may immediately observe that for every component of other than For every such where there exists a vertex we set For all other components we set Moreover, for each component of other than we set
Notice that we have and in all cases. Hence, we may now apply our induction hypothesis to every pair and This either yields a -segregated grid as a colorful minor of or for each we find a rooted tree-decomposition satisfying the properties required by the assertion and with
Let now be the tree obtained from the disjoint union of all by introducing a new vertex a new vertex and then joining with an edge to and all Finally, we set and for every and It follows directly from our construction that is a rooted tree-decomposition precisely as required by the assertion. ∎
6 The Erdős-Pósa property for -colorful graphs
With our structural theorems in place, we are now ready to proceed with the proof of theorem˜1.5. To this end, we first define, for every the sets and (section˜6.1); we then introduce the class of crucial colorful graphs (section˜6.1). The remainder of the section is devoted to proving that a -colorful graph has the Erdős-Pósa property if and only if it is crucial (theorem˜6.19). In the proof, the set serves as the (colorful minor) obstruction set for the class of all -colorful graphs that satisfy the Erdős-Pósa property. The “upper bound” part of the proof is presented in section˜6.3 (connected case) and section˜6.4 (non-connected case), while the “lower bound” is established in section˜6.5.
6.1 The sets and
In this section we define, for every five sets of -colorful graphs which we denote by and We prove that they determine precisely which colorful graphs have the Erdős-Pósa property.
Planarity.
We start by defining the set where assigns to each vertex the empty set for every The following is the Kuratowski-Wagner characterization of planar graphs [88].
Observation 6.1.
For every non-negative integer a -colorful graph is planar if and only if it does not contain a member of as a colorful minor.
Color facial graphs.
With we have a way to control the global structure of our graphs. Next, we begin to further restrict the structure with respect to the colors. We say that a colorful graph is color-facial if it has a plane embedding where all colored vertices are on the boundary of a single face.
Observation 6.2.
For every non-negative integer the class of color-facial -colorful graphs is closed under colorful minors.
We denote by (resp. ) the graph obtained from (resp. ) by deleting a single edge. For each we define to be the set of all the -colorful graphs and where for every and each
assigns one or two colors from to the two vertices of that have degree and no vertex receives two colors at once – all other vertices are assigned
assigns one, two, three, or four colors from to the vertices of such that every vertex receives exactly one color,
assigns one or two colors from to the two vertices of of degree such that no vertex receives colors at once – all other vertices are assigned and
assigns one, two, or three colors from to the three vertices of of degree such that no vertex receives colors at once – all other vertices are assigned
We set
Observation 6.3.
For every non-negative integer the family is an anti-chain for the colorful minor relation.
Lemma 6.4.
A -colorful graph is color-facial if and only if it does not contain a member of as a colorful minor.
Proof.
Let be the graph obtained from by we adding a new vertex and making it adjacent to all vertices with
“” Let be color-facial. Since the property of being color-facial is closed under colorful minors, every colorful minor of should also be color-facial. Notice that none of the colorful graphs in are color-facial, therefore all of them are excluded from as colorful minors.
“” Assume now that is not color-facial. Then, is not planar, since otherwise, by considering a plane embedding of and removing all (colored) neighbors of are in the boundary of a face, i.e., is color-facial, a contradiction. As is not planar it contains a subgraph that is either a subdivision or a subdivision of We call a vertex of a branch vertex if it has degree at least in otherwise we call it a subdivision vertex.
If is not a vertex of then should be a subgraph of as well, therefore contains or as a minor, which means that contains some of the graphs in as a colorful minor. Therefore we may assume that and distinguish cases:
If is a subdivision of and is a subdivision (resp. branch) vertex. Then all (resp. ) neighbors of in have a non-empty palette. Therefore, must contain some (resp. ) as a colorful minor.
If is a subdivision of and is a subdivision (resp. branch) vertex. Then, all (resp. ) neighbors of in have a non-empty palette, therefore contains some (resp ) as a colorful minor. ∎
We define by discarding all -colorful graphs with and every whose vertex set can be partitioned into two sets of size such that there are and For example, in the case where the only graph discarded is the one in the dashed square in figure˜5.
Notice that contains variants of the graphs and each. Moreover, and give rise to a total of colorful graphs each. In total, for every
Color-segmentation.
We say that a -colorful graph is color-segmented if the three following conditions are satisfied:
A.
for every cycle with four distinct vertices appearing in this order on there are no colors such that
B.
every cycle contains at most two vertices carrying more than one color, and
C.
for every vertex among the connected components of that contain neighbors of there are at most two that carry more than one color.
For let be the set containing
every -colorful graph where each vertex carries exactly one color and such that, for any two non-adjacent vertices, their palettes are disjoint from those carried by the other two.
every -colorful graph where for each
every -colorful graph where for each leaf of
If then See figure˜6 for an illustration.
By the definitions, it is easy to observe the following.
Observation 6.5.
For every non-negative integer the class of color-segmented -colorful graphs is closed under colorful minors.
Observation 6.6.
For every non-negative integer is an anti-chain for the colorful minor relation.
Observation 6.7.
For every non-negative integer a -colorful graph is color-segmented if and only if it does not contain any member of as a colorful minor.
In this case, we define to be the set obtained from by discarding all with
It follows that each of the basic types, contributes a total of -colorful graphs to the family Therefore,
Excluding three colors from a component.
We say that a -colorful graph is component-wise bicolored if for every component of
For we define as the set containing every -colorful graph where assigns out of the colors to the unique vertex of For instance, a visualization of is
For we set Clearly, for it holds that The next observation is straightforward.
Observation 6.8.
For every non-negative integer the class of component-wise bicolored -colorful graphs is closed under colorful minors.
Observation 6.9.
For every non-negative integer the family is an anti-chain for the colorful minor relation.
Observation 6.10.
For every non-negative integer a -colorful graph is component-wise bicolored if and only if it does not contain any member of as a colorful minor.
Further restricting the colors of the components.
For every non-negative integer we call a -colorful graph multicolored if
We say that a -colorful graph is single-component bicolored if for any two components and of there are no four distinct colors such that and
For we define to be the set containing every -colorful graph where is such that we have for both and
As an example, consider the set
Observe that for every non-negative integer it holds that
The next three observations are immediate.
Observation 6.11.
For every non-negative integer the class of single-component bicolored -colorful graphs is closed under colorful minors.
Observation 6.12.
For every non-negative integer the family is an anti-chain under the colorful minor relation.
Observation 6.13.
A -colorful graph is single-component bicolored if and only if does not contain any member of as a colorful minors.
6.2 Crucial colorful graphs
Let be a -colorful graph. We say that is crucial if the following conditions are satisfied:
i)
is color facial,
ii)
is color segmented,
iii)
is single-component bicolored, and
iv)
is component-wise bicolored.
Recalling Observations˜6.2, 6.5, 6.8 and 6.11, one can see immediately that crucial colorful graphs are also closed under taking colorful minors.
Observation 6.14.
For every non-negative integer the class of crucial -colorful graphs is closed under colorful minors.
We now need to justify why we excluded some colorful graphs from in order to define for
Lemma 6.15.
For every non-negative integer the family contains exactly the members of that do not contain some member of as a colorful minor. Moreover, contains exactly the members of that do not contain any member of as a colorful minor.
Proof.
By the definition of if a -colorful graph is not contained in as a colorful minor, this is because either it is not component-wise bicolored or it is where two vertices carry the color and the other carry a different color In the first case, contains a colorful graph from as a colorful minor, while in the second, the -colorful graph contains the -colorful graph as a colorful minor where has two non-adjacent vertices satisfy and the other two, say satisfy Notice now that none of the -colorful graphs in contain any of the -colorful graphs in as a colorful minor. This is because every cycle of carries at most two colors and one of those colors appears at most once. Also as the same colorful graphs carry two colors, they cannot contain any of the colorful graphs in as colorful minors.
For the second statement, recall that in the definition of we discarded all members of carrying more than two colors and these are exactly those that contain some member of as a colorful minor. ∎
For each non-negative integer we now define
Taking into account lemma˜6.15, it is easy to verify the following.
Observation 6.16.
For every non-negative integer the family is an anti-chain for the colorful minor relation.
Lemma 6.17.
For every non-negative integer a -colorful graph is crucial if and only if it does not contain any member of as a colorful minor.
Proof.
By observation˜6.16 we know that no member of contains another as a colorful minor. Moreover, by lemma˜6.4 we have that no member of is color-facial, no member of is color-segmented by observation˜6.7, no member of is single-component bicolored by observation˜6.10, and no member of is component-wise bicolored by observation˜6.10. Hence, no crucial -colorful graph can contain any member of as a colorful minor.
For the reverse direction let be a -colorful graph that does not contain any member of as a colorful minor. Due to lemma˜6.15 and by lemmata˜6.4 and 6.7 we know that is color-facial and color-segmented. Moreover, by observation˜6.10 we have that is single-component bicolored and by observation˜6.13 it is also component-wise bicolored. Hence, is crucial. ∎
We are also able to obtain the size of precisely. Our previous discussions on the families and imply that
Interestingly, these numbers, being polynomially bounded, are rather tame. Indeed, for some small values of we get and
Packings and half-integral packings.
Let be a non-negative integer and let and be two -colorful graphs. Fix a positive integer A packing of in of size is a collection of pairwise vertex-disjoint -colorful graphs such that is a subgraph of and is a colorful minor of for each
A half-integral packing of in of size is a collection
of -colorful graphs such that is a subgraph of and is a colorful minor of for each and no vertex of appears in more than two of the
It is not hard to see that for any permutation of the -segregated grid realizing on its set of colors contains an half-integral packing of size of the -segregated grid which also realizes on its set of colors. See figure˜7 for an example.
Observation 6.18.
For all positive integers and every permutation of the -segregated grid realizing contains a half-integral packing of size of the -segregated grid realizing
Erdős-Pósa property for colorful graphs.
Let be a non-negative integer and be a -colorful graph. We say that has the Erdős-Pósa property if there is a function such that for every and every -colorful graph either there is a packing of of size in or there is a vertex set such that and does not contain as a colorful minor. We call the function the gap of the Erdős-Pósa property of
We may now state the main result of this section.
Theorem 6.19.
For every non-negative integer a -colorful graph has the Erdős-Pósa property if and only if it is crucial.
The proof is divided into three parts spanning the following three subsections. First we show that every connected and crucial -colorful graph has the Erdős-Pósa property. Then we prove that this can be extended to cover also all non-connected crucial -colorful graphs, and finally, we show that -colorful graphs that are not crucial do not have the Erdős-Pósa property.
6.3 The positive part of theorem˜6.19: The connected case
We prove here that if is crucial and connected then it has the Erdős-Pósa property.
Let us begin with a small observation. For there exist permutations such that there exists some for which the -segregated grid realizing is not a colorful minor of the -segregated grid realizing for any possible choice for See figure˜2 for three such segregated grids in the case where However, for there exists only one segregated grid over all. The case is also special since there exist two cyclic permutations on however, it is easily seen that both permutations give rise to the same family of segregated grids up to isomorphism.101010In general, for every positive integer there are different cyclic permutations without direction. Moreover, there exists a non-negative integer such that every -segregated grid is contained as a colorful minor in the -segregated grid corresponding to one of these permutations, for every non-negative integer
The following is a simple observation on the behavior of grids. See figure˜8 for an illustration.
Observation 6.20.
For each and every choice of positive integers and The -segregated grid contains a packing of size of the -segregated grid.
The following is a simple observation based on the proof of Robertson, Seymour, and Thomas [80] that every planar graph is a minor of some grid that takes into account that is color facial and color segmented.
Proposition 6.21.
For each there exists a function such that every crucial -colorful graph is a colorful minor of the -segregated grid.
Combining proposition˜6.21 with observation˜6.20 immediately yields that for large -segregated grids yield large packings for any crucial -colorful graph. Indeed, this observation can easily be extended as follows.
Observation 6.22.
For every integer there exists a function such that every connected crucial graph is a colorful minor of some -segregated grid.
The above property is implied by crucial graphs being single-component bicolored. Finally, we also need to observe how being component-wise bicolored implies a similar property. See figure˜9 for an illustration.
Observation 6.23.
For every integer there exists a function such that for every positive integer and every every -segregated grid contains vertex-disjoint subgraphs such that for every it holds that and contains a -segregated grid as a colorful minor.
Corollary 6.24.
For every non-negative integer and every crucial -colorful graph there exists a function such that for every positive integer the -segregated grid contains a packing of size of
Observe that corollary˜6.24 combined with theorem˜1.3 directly implies the following.
Corollary 6.25.
For each non-negative integer and every crucial -colorful graph there exists a function such that for every non-negative integer and every -colorful graph either
i)
contains a packing of size of or
ii)
the torso treewidth of is at most
Lemma 6.26.
For every non-negative integer all non-negative integers every -colorful graph and every -colorful graph together with a set such that each component of is restricted and where the treewidth of the torso of in is at most one of the following is true:
i)
There is a set of size at most such that no component of contains as a colorful minor,
ii)
and there is a set of size at most such that every component of is either restricted or does not contains as a colorful minor, or
iii)
there is a packing of in of size
Proof.
For any -colorful graph let be the maximum size of a packing of in and be the minimum size of a set such that no connected component of satisfies and contains as a colorful minor. We show that We prove this by induction on
By assumption there is a set such that and for every component of the -colorful graph is restricted. This implies that there is a tree decomposition of with adhesion at most and with a special set of leaves such that for every node and where for every leaf node adjacent to node is restricted. Given an edge of we denote, for each by the component of that contains Moreover, we define
Now, assume there is an edge such that and Then we have that and for both By calling upon the induction hypothesis it follows that
Hence, in this case there is nothing more for us to show and we may assume that for any edge there exists such that
Let us now further assume that there exists some edge such that and This implies that no component of contains as a colorful minor and since our proof is complete.
The only remaining case is where for every edge of there exists precisely one such that In this case, we orient each towards the component of where the value of is positive. This means that there is a node of which is a sink of the resulting orientation of
Let us assume that is not adjacent to any member of It follows that We now have that no connected component of contains as a colorful minor and as as claimed.
Thus, we may now assume that is adjacent to some member of Let be the neighbors of in Consider, for every the edge and assume that there exists some such that and Notice that, similar to before, we have Let us apply our induction hypothesis to If this returns a packing of size we are done. Hence, we may assume and thus, by the induction hypothesis, we have that Moreover, we have that is restricted and Hence, it follows that
Therefore, we may now assume that for every either or or both. If both are equal to for some we enter a similar case as before since now deleting ensures that none of the remaining components can carry all colors and contain as a colorful minor. Thus, we may assume that for each precisely one of and is non-zero. As before, this now allows us to orient each edge towards the side where is non-zero. If, in the resulting orientation, has no outgoing edge, this means that every component of either is restricted or does not contain as a colorful minor as desired. Hence, we may assume that has at least one outgoing edge. Indeed, this implies that has a unique outgoing edge, say by our assumption. In this case, however, we know that and, by the definition of and our orientation, we have that every component of is either restricted or does not contain as a colorful minor.
Finally, notice that if is a colorful minor of some restricted subgraph of then this implies that as required by the second outcome or our assertion. ∎
We first utilize lemma˜6.26 to show that connected crucial colorful graphs have the Erdős-Pósa property. For this, recall that planar graphs have the Erdős-Pósa property for minors due to a result of Robertson and Seymour [76].
Proposition 6.27 (Robertson and Seymour [76]).
A (-colorful) graph has the Erdős-Pósa property if and only if is planar.
Theorem 6.28.
For every non-negative integer every connected and crucial -colorful graph has the Erdős-Pósa property with gap where is the gap from proposition˜6.27.
Proof.
For the assertion follows directly from the classic result of Robertson and Seymour, i.e. proposition˜6.27. Hence, we may assume
Fix any crucial -colorful graph and consider an arbitrary -colorful graph Next, let us adjust as follows. For each let We also set Then both and are -colorful graphs.
We first call upon corollary˜6.25. This allows us to assume that the torso treewidth of is at most as otherwise, corollary˜6.25 would immediately provide us with a packing of size of within – and therefore within
Hence, by applying lemma˜6.26 we now either find a packing of size of in or there is a set of size at most such that no component of carries all colors and contains as a colorful minor. However, by our choice of and as well as the assumption that is connected, the above implies that no restricted subgraph of may contain as a colorful minor. Moreover, since none of the colors in have any influence on whether is a colorful minor of or not, it follows that also does not contain as a colorful minor and our proof is complete. ∎
6.4 The positive part of theorem˜6.19: The non-connected case
We continue with the proof of theorem˜6.19, showing how we can reduce the case of general crucial -colorful graphs to theorem˜6.28 where is connected.
Theorem 6.29.
For every non-negative integer and every crucial -colorful graph there exists a function such that has the Erdős-Pósa property with gap
Proof.
Let denote the number of connected components of and let be the set of non-empty proper subsets of We define a recursive function as follows:
The function is a placeholder for the actual function we wish to define to prove the assertion. However at the moment it is more convenient to make the dependency on the number of connected components of visible. To obtain the result we wish for, we may simply set as is always an upper bound on
We prove the following by induction on the number of connected components of
Inductive statement: For every non-negative integer and all positive integers and every crucial -colorful graph with at most components has the Erdős-Pósa property with gap
Base of the induction: The base follows in a straightforward manner from theorem˜6.28. This is true since for the base we may consider the case where and is connected.
Inductive step: From here on, we may assume that Let be the components of For every we define the -colorful graph
Consider a -colorful graph and set for all Then, assuming and are both -colorful graphs and for any subgraph is a colorful minor of if and only if is a colorful minor of Observe that either there is a packing of in of size or by calling upon corollary˜6.25, we may conclude with a set such that and every component is restricted. By lemma˜6.26 we conclude with a set of size at most such that no connected component of contains as a colorful minor.
This implies that for every minimal colorful subgraph of that contains as a colorful minor there exists a partition of such that where for every is a minimal colorful subgraph of a connected component of that contains as a colorful minor, and each is a subgraph of a different connected component of than all others. Let us call this the partition property.
Let For every let be the set of all connected components of that contain as a colorful minor. Moreover, for any set of connected components of we define as the set of all connected components of not in We first distinguish two cases, one of which is easily dealt with, and then we enter, as part of the second case, an iterative construction process that eventually leads to either the desired packing or the desired set.
Case 1: for and
Case 2: there exists such that
Notice that, in Case 1, we may greedily choose a partition of and select components from and thereby obtain a packing of in of size Therefore, we may assume to be in Case 2. In the following, we will, at times, apply a similar counting argument. To facilitate our writing, we will simply refer to this as the few components argument.
We initialize and iteratively construct sets and as follows:
Target of our construction: For every there exists a set such that there exists a set of size at most
and a set of at most connected components of with the following properties:
i)
for every and every connected component either contains a packing of of size or does not contain as a colorful minor, and
ii)
for every and every connected component it holds that does not contain as a colorful minor.
Once we complete this construction, we obtain that During this construction, we stop whenever we reach the point where does not contain as a colorful minor, or contains a packing of of size For the second outcome of this stopping condition we will employ a small argument based on Hall’s Theorem. This argument will also complete the final case
We proceed with the construction by induction on and, as the base, show how to obtain the set
Base of our construction: We set Notice that the second condition holds by the definition of since no other component of can contain as a colorful minor. For each of the at most members of we apply the induction hypothesis on Let be the union of all hitting sets of size at most returned in this way. It follows that and also satisfy the first condition of our construction.
The inductive step of our construction: Let us assume that for some we have already constructed the sets and as above.
First, let us introduce the matching argument. Consider any partition of such that is non-empty. Let be the bipartite graph with as one of the two color classes and as the other color class. There exists an edge if and only if for the component still contains as a colorful minor. Notice that, by the second condition of our construction, this means that must contain a packing of of size If does not have a matching that covers the entire set then, by Hall’s Theorem, there exists a set such that This implies that no colorful minor of can be realized in via the partition Now, by the partition property, if for every partition of is non-empty and does not contain a matching that covers then does not contain as a colorful minor and we may terminate. Hence, we may assume that there is a partition of for which either is empty or has a matching that covers In the latter case, this implies that has a packing of of size
Now, suppose that for By the few components argument and the matching argument applied to we conclude that contains a packing of of size in any of the two cases above. Hence, we may assume that there is such that Let
Next, we apply the induction hypothesis on and every member of Let be the union of all hitting sets, each of size at most which are returned by these applications. Notice that Finally, let and notice that
The three sets and now satisfy the first condition of our construction by the induction hypothesis and our construction of and the second condition is satisfied by our choice of
In particular, it follows from the matching argument that, if for every partition of such that is non-empty, does not contain as a colorful minor, then no colorful minor of can be realized in via the partition This last observation indeed proves that, once and are constructed, we are done. ∎
6.5 Negative part of theorem˜6.19
In this section we show that if is not crucial then it does not have the Erdős-Pósa property. This is the last remaining part for the proof of theorem˜6.19.
Let be a non-negative integer and and be -colorful graphs. We denote by the smallest integer such that there exists a set with and does not contain as a colorful minor. Similarly, we denote by the largest integer such that there exists a packing of of size in
In this terminology, we may rephrase our definition of the Erdős-Pósa property as follows: A -colorful graph has the Erdős-Pósa property if there exists a function such that for every -colorful graph it holds that
We begin with a simple but important observation.
Lemma 6.30.
For every non-negative integer the class of -colorful graphs that have the Erdős-Pósa property is closed under colorful minors.
Proof.
Let and be -colorful graphs where is a colorful minor of and assume that has the Erdős-Pósa property. This means that there is a function such that for every -colorful graph Notice that, as is a colorful minor of
We conclude that, for every -colorful graph
as required. ∎
The following is a known fact about the embeddability of graphs in surfaces of positive Euler genus.
Proposition 6.31 (see e.g. [63]).
Let be two graphs, each containing or as a minor, and let be the result of the identification of some vertex of and some vertex of Then is not embeddable in the projective plane.
Theorem 6.32.
For any non-negative integer if a -colorful graph has the Erdős-Pósa property, then is crucial.
Proof.
Due to lemmata˜6.30 and 6.17, it is enough to prove that none of the colorful graphs in has the Erdős-Pósa property. We first examine the case where is a connected -colorful graph from Clearly cannot be a member of as proposition˜6.27 already covers all non-planar graphs.
Before we proceed with the proof we present a construction together with a preliminary observation, in the form of Claim 0 as stated below.
Let be a -colorful graph and let be a single-boundaried surface (i.e., a surface from which we removed an open disk) where is embedded in a way that all vertices that carry colors are drawn on the boundary of The -multiplication of in is the colorful graph created by taking drawings of in chosen so that no point of is in more than two of them, only finitely many points of are in more than one of the copies, and any two distinct copies only intersect in edges. Moreover, we insist that no crossing point is drawn on the boundary of and that each vertex of that was drawn on the boundary of gives rise to new vertices of which are now drawn consecutively on the boundary of We may now replace each crossing by a new vertex of degree that does not carry any color and declare the resulting -colorful graph to be where every vertex from one of the copies of inherits its palette.
Claim 0: For every if is the -multiplication of in then
Notice that has subgraphs each a subdivision of a copy of such that no vertex of is in more than two of the Moreover, is isomorphic to for each Thus, fewer than vertices cannot cover all of those therefore
Assume now that is a -colorful graph that has the Erdős-Pósa property. Our aim is to prove a series of lemmas that, one by one, will establish the properties for to be crucial.
Claim 1: is not a member of
Assume the contrary and let be the graph obtained from by adding a new vertex and making it adjacent to all vertices of that carry colors under Clearly is either or a subdivision of one of them. Consider an embedding of in the projective plane As all colored vertices of are neighbors of we obtain that can be embedded in in a way that all colored vertices are on the boundary of a disk of We now choose to be the surface obtained from by deleting the interior of We also define as the graph obtained from by introducing a new vertex and making in adjacent to all vertices of which are drawn on the boundary of Clearly can be drawn in
It is enough to prove that as this, due to Claim 0, would contradict our assumption that has the Erdős-Pósa property.
Assume to the contrary that contains two disjoint subgraphs and such that is a colorful minor of for both For let be the graph obtained by adding the vertex and joining it to all vertices of with a non-empty palette under Clearly both and are subgraphs of and their only common vertex is Notice that each contains one of as a minor. By proposition˜6.31 this means that must contain a minor that is not embeddable in the projective plane which is absurd.
So far, we have ruled out the cases where Given that is color-facial we know that has an embedding in a disk where all vertices of with a non-empty palette are drawn on the boundary of Hence, we may assume that This means, in particular, that either or has at least one vertex whose palette contains more than one color.
We now “uncontract” each vertex where to a path on whose vertices are drawn on the boundary of in a way that each vertex of this path carries exactly one of the colors of This creates a new -colorful graph which contains as a colorful minor and, moreover, each colored vertex has exactly one color.
By its construction, is also embedded in such that all colored vertices are drawn on its boundary. For each let now denote the -multiplication of in Notice that, in the construction of each vertex of gives rise to vertices appearing consecutively on the boundary of and all carrying the same color under We call each such sequence of vertices a color interval of Clearly, as is a -colorful graph, has at most color intervals. Indeed, since by definition of we have that has at most color intervals.
By applying Claim 0 to and we obtain that As is a colorful minor of we have that This implies that
Let and assume that contains pairwise-disjoint connected subgraphs such that is a colorful minor of for each We may further assume that the ’s are subgraph minimal. Recall that, by the construction of each vertex of caries at most one color for each
Claim 2: is not a member of
Assume the contrary and observe that, by minimality, each is the union of three (possibly trivial) paths with a common endpoint and such that the other three endpoints have distinct colors. By identifying each of the color intervals in the boundary of we can now see the vertices resulting from this identification as the vertices of a hypergraph whose hyperedges correspond to the subgraphs of have arity and can be drawn as pairwise-disjoint open disks in the interior of a closed disk such that their endpoints are on the boundary of A simple counting argument implies that which, along with (1), implies that does not have the Erdős-Pósa property, a contradiction.
Claim 3: is not
Assume the contrary and observe that, by minimality, each is the union of a cycle and four (possibly trivial) paths with one endpoint on and the other carrying one of two colors in a way that appear in this order in the cycle and carry one color and carry a different one. By identifying each of the color intervals on the boundary of we can now see the subgraphs as hyperedges of arity drawn as pairwise-disjoint open disks in the interior of a closed disk whose endpoints are on the boundary of and the total number of their endpoints is at most the number of color intervals in the boundary of Hence, there are at most such endpoints. A simple counting argument implies that which, as in the previous claim, contradicts the Erdős-Pósa property for
Claim 4: is neither or
Proof of Claim 4: As before we assume the contrary. Observe that the minimality of the implies that – depending on the case – each can be constructed by considering the disjoint union of three disjoint paths each containing two distinct vertices colored by two distinct colors and and by adding three new and pairwise internally-vertex-disjoint paths and which are also internally-vertex-disjoint from as follows:
for all one endpoint of is in the other is in and
have a common endpoint that does not belong to while, for each the other endpoint of is in where
We now claim that, in both cases, it is impossible for two color intervals, one of color and one of color to host the endpoints of and To see this we assume that there exists a where all of the only interact with two color intervals and construct an auxiliary graph as follows: In case the of we add a new vertex and make it adjacent to both endpoints of each path Moreover, we add a cycle along the boundary of (see the left side of figure˜10). In the case the of we add a new vertex and make it adjacent to the endpoints of the colored by We then introduce a second new vertex and make it adjacent to the endpoints of the colored by (see the right side of figure˜10). In both cases, the resulting graph has an embedding in the plane, by drawing precisely the new vertices in the complement of with respect to the plane. However this not possible as, in any case, the resulting graph, by construction, contains either or as a minor. We conclude that each occupies vertices of at least three color intervals. Therefore, as before, we can identify each of the, at most color intervals in the boundary of and see the graphs as hyperedges of arity drawn as open disks in the interior of a closed disk. Similar to previous claims, we may now observe that As before, this is a contradiction to the Erdős-Pósa property of
We just proved that if is connected and has the Erdős-Pósa property, then it cannot be a member of It remains to prove that does not have the Erdős-Pósa property.
Let be a -colorful graph that is the disjoint union of two (non-trivial) paths and whose endpoints are colored 1, 3, and 2, 4 respectively and whose internal vertices have empty palettes. Consider an embedding of in the closed disk with one crossing so that the colored vertices are drawn on the boundary of We next consider the -multiplication of in as indicated in figure˜11. The resulting graph has a drawing in with crossings where every vertex carries at most one color. Notice also that all the crossing points are homotopic in
With arguments analogous to those we used in the proof of Claim 0 it easily follows that Consider two disjoint paths in say and with endpoints colored and respectively. Observe that, in all components carry at most two colors. Moreover, those colors are consecutive in the cyclic ordering We conclude that which proves that does not have the Erdős-Pósa property and completes our proof. ∎
The two directions of theorem˜6.19 now follow from theorem˜6.29 and theorem˜6.32.
As a special case we investigate the Erdős-Pósa property for -colorful graphs that are rainbow, i.e. the palette of every vertex is As a consequence of theorem˜6.19 we obtain the following.
Theorem 6.33.
Let be a non-negative integer and be a -colorful rainbow graph. Then has the Erdős-Pósa property if and only if and excludes and as minors. In other words:
if then is planar,
if then is outerplanar, and
if then is a union of paths.
Proof.
Case is obvious as this is precisely proposition˜6.27.
For case recall that a graph is outerplanar if and only if it excludes both and as minos. Observe that if a graph contains or as a minor then the -colorful rainbow graph contains one of the -colorful graphs or from as a colorful minor. This, by theorem˜6.19, it would not have the Erdős-Pósa property. For the inverse, because the graph of every -colorful graph from contains either or as a minor, for every -colorful graph that contains a member of as a minor it holds that must also contain or as a minor (see figure˜5).
For the case recall that a graph is a disjoint union of paths if and only if it excludes both and as minors. Observe that if a graph contains or as a minor then the -colorful rainbow graph contains one of the -colorful graphs or from as a colorful minor. This is because every -colorful graph in contains either or as a minor. This also implies that if contains some graph from as a colorful minor then also contains or as a minor.
The case is impossible because then some of the -colorful graphs in is contained as a colorful minor in . ∎
theorem˜6.19 and observation˜6.16 imply that, for every is the colorful minor obstruction set of the class of -colorful graphs that have the Erdős Pósa property.
Given some we define the class as follows: contains the -grid without any colors, contains the -segregated grid and, for contains every -colorful graph that is the disjoint union of the set of all non-isomorphic -segregated grids with some common color (see figure˜12). From observation˜6.22 and observation˜6.23 it follows that every crucial -colorful graph is a colorful minor of some graph in for some Notice that, while for as there are exactly ways to choose the common color. Moreover it is direct to verify that every -colorful graph in is crucial. This permits us to restate theorem˜6.19 as follows.
Corollary 6.34.
Let be a non-negative integer. The set is the obstruction set of the class of all -colorful minors of the -colorful graphs in This class is precisely the class of all -colorful graphs that have the Erdős Pósa property.
7 Algorithmic applications
To present the algorithmic applications of our results, we first introduce a unifying concept that encompasses all of them.
A graph parameter is called decomposable if it is subgraph-monotone and, for every separation of we have Examples of decomposable parameters considered in this paper include treewidth and the Hadwiger number
7.1 Star Decompositions of Colorful Graphs
Given a tree with at least one vertex, we denote by the set of its leaves. A star is any connected subgraph of for some non-negative integer If then has exactly one non-leaf vertex, called its center. If any vertex of may be designated as the center.
A star decomposition of a -colorful graph is a tree decomposition where is a star with center and i.e. all vertices with non-empty palettes appear in the center’s bag. Without loss of generality, we assume is chosen so that is maximized. This ensures that for each the set is connected in
For a graph parameter the -width of is defined as
The following observation follows by iteratively applying the definition of parameter decomposability to each leaf in a star decomposition.
Observation 7.1.
Let be a decomposable parameter. If a graph admits a star decomposition of -width at most then
The next two results follow from theorem˜5.3 and theorem˜3.6, respectively, applied to -colorful graphs. These structural results can be stated as follows:
Corollary 7.2.
There exists a function and an algorithm that, given a -colorful graph and an integer either outputs a colorful minor model of a -segregated grid or a star decomposition of with -width at most in time.
Corollary 7.3.
There exists a function and an algorithm that, given a -colorful graph and an integer either outputs a colorful minor model of a -colorful rainbow or a star decomposition of with -width at most in time.
In the following sections, we present algorithmic applications of corollary˜7.2 and corollary˜7.3.
7.2 An algorithmic reduction
We now introduce the main algorithmic machinery of this section. It consists of an algorithmic reduction that relies on two key algorithmic assumptions and the existence of a star decomposition along with a graph parameter associated with it. The first assumption is that, for the problem under consideration, there exists an efficient method for replacing boundaried subgraphs of a colorful graph with equivalent but smaller ones. The second assumption is that the problem can be solved efficiently when the value of is bounded.
Equivalences and representations.
For every non-negative integer we denote the class of all -colorful graphs by We define a functional problem111111We use the term “functional problem on -colorful graphs” instead of “-colorful graph parameter” to emphasize that we are addressing a computational problem. on -colorful graphs to be a map Our goal is to identify conditions that enable the design of algorithms for solving functional problems on -colorful graphs.
We say that two boundaried -colorful graphs and are -equivalent if they are compatible and, for every boundaried -colorful graph that is compatible with (and therefore with as well), it holds that
where given two compatible boundaried -colorful graphs and denotes the boundaried -colorful graph obtained by identifying the vertex with the vertex for every
A boundaried -colorful graph is confined if A representation of is a pair where maps confined boundaried -colorful graphs to -equivalent confined boundaried -colorful graphs, and maps integers to integers such that, for every confined boundaried -colorful graph the graph has at most vertices. We emphasize that the notion of confinedness plays a crucial role in both our definitions and results.
Theorem 7.4.
Let be a functional problem on -colorful graphs, and be a decomposable graph parameter. Suppose that there exists functions such that the following hold:
i)
There exists an algorithm that, given a -colorful graph computes in time.
ii)
There exists a representation of and algorithms that, given a boundaried -colorful graph computes and in and time respectively.
Then, there exists an algorithm that, given a star decomposition of a -colorful graph with -width at most outputs in time.
Proof.
Let be the center of For each leaf of consider the boundaried -colorful graph where is a bijection from to Notice that is confined by definition of a star decomposition. Using the algorithm from (ii), compute Initialize For each leaf update
effectively replacing with in By -equivalence it holds that This construction runs in time.
Now consider the star decomposition of where is the vertex set of for each By observation˜7.1 we know that Using the algorithm from (i), we compute (which is equal to ) in time, as required. ∎
All subsequent meta-algorithmic results will rely on theorem˜7.4.
7.3 Logics and AMTs on colorful graphs
We now present the logical frameworks for two Algorithmic Meta-Theorems (AMTs) that we will extend to the colorful setting.
CMSOLogic on Colorful Graphs.
We consider logical formulas on colorful graphs with the following variables: A vertex set variable and an edge set variable Also we consider a collection of set variables representing colors. A -colorful graph is represented as the structure where: interprets interprets and for interprets These variables are always free in We also consider: Sequences of vertex/edge variables / and vertex/edge set variables / which are quantified in formulas. These formulas describe structures where are indexed sets interpreting the corresponding variables. For instance, for some where interprets We call the index of with analogous definitions for and
A monadic second-order logic (MSO) formula may be:
An atomic formula:
“” (vertex equality),
“” (vertex adjacency),
“” or “” (membership).
A compound formula built using and (quantifying over vertex/edge/set variables).
For a structure and an index not in ’s index, we say if there exists a vertex such that models with interpreted as Quantifiers and are defined similarly. We write as shorthand for with analogous definitions for and
A counting monadic second-order logic (CMSO) formula extends MSO with predicates which hold when is divisible by For a colorful graph and CMSO-formula we write if models
The Logic.
The logic modifies by replacing: with () and with We next define the special quantifier : For a structure and index not in ’s index, if there exists a vertex set such that and, moreover, The quantifier is defined analogously, requiring for the edge set ’s endpoints 121212In the definition of [81] it is demanded that that is the maximum treewidth of an -rooted minor of Clearly, and can be seen as -colorful graph parameters that are equivalent because of the main result in [12]. In this paper we prefer to define using instead of as this is more adequate to the colorful setting of the paper. Also, given a formula we use the notation in order to denote the bidimensionality bound that it imposes on the interpretation of In order to unify notation, in case of a CMSO-formula where there is no restriction on quantification, we assume that
Disjoint Paths Extensions.
Given a we next define the -ary disjoint-paths predicate The logic extends by allowing atomic formulas of this form. For a structure holds if there exist vertex-disjoint paths in connecting each to
We define CMSO logic as the set of all CMSO-formulas and, similarly, we define logic. We stress that has the same expressibility as as the disjoint paths predicate can be expressed by -formulas. However, is a proper subset of which is more expressive as the disjoint paths predicate is not expressible in
CMSO-Definable Functional Problems on -Colorful Graphs.
Let be a non-negative integer, and let be an CMSO-formula over -colorful graphs. We define the maximization problem on as the functional problem where
where i.e., we add a new color to all vertices of We also define the minimization problem analogously, by replacing “” with “” and “” by “” in the above.
In general, when we do not need to specify we simply use the term optimization problem (on ). Notice that in (2) we insist that the vertex set carrying the “new color” is a subset of the set of vertices that already carry some color.
The following result is the optimization extension of Courcelle’s Theorem ([15, 13]) proved by Arnborg and Andrzej Proskurowski [4], stated in terms of colorful graphs (see also [9, 3]).
Proposition 7.5.
There exists an algorithm that, given a formula over -colorful graphs, and a -colorful graph computes in time
Sau, Stamoulis, and Thilikos [81] proved the following Algorithmic Meta-Theorem (AMT). While in its full generality the result is stated for general structures, we present it here for colorful graphs. Actually, we present the following optimization version of it, that can be easily derived from the results of [81] (for this optimization version, it is important to demand in the definition of in (2)).
Proposition 7.6.
There exists an algorithm that, given a formula over -colorful graphs, and a -colorful graph computes in time
We say that a -colorful graph parameter is CMSO-definable by some MSO-formula on -colorful graphs and some choice of if If is a -formula, then we say that is -definable (by and ).
By proposition˜7.5 (resp. proposition˜7.6), CMSO-definability (resp. -definability) ensures Condition (i) of theorem˜7.4 when (resp. ). This naturally raises the question of what assumptions may guarantee Condition (ii). One answer, explored in this section, involves the notion of topological minor folios.
7.4 Folio representation
theorem˜7.4 has two algorithmic conditions. The first one will be derived from the conditions of propositions˜7.5 and 7.6. The second will follow from the ability to replace some part of the input colorful graph with another in which we can route the same set of rooted topological minors (up to a given size, depending on the problem). Provided that this replacement can be performed algorithmically, it will yield an additional combinatorial condition for our algorithmic meta-theorems.
Topological folios.
Especially for boundaried -colorful graphs, we omit the empty coloring and simply write triples which we call boundaried graphs.
Given two boundaried graphs and we say is a topological minor of if there exists a pair where:
is a subgraph of with and all vertices in have degree 2,
Dissolving all vertices in yields a boundaried graph isomorphic to
Notice that if is a topological minor of and then and are compatible confined boundaried -colorful graphs.
The -topological-folio of a confined boundaried -colorful graph is the set of all confined boundaried -colorful graphs where is a topological minor of and We denote this set by
Theorem 7.7.
There exists a computable function such that for every and every confined boundaried -colorful graph there exists a (compatible) confined boundaried -colorful graph where: and
Proof.
This follows from [42, Lemma 2.2] for boundaried graphs. The confined -colorful case holds because depends only on not on ∎
theorem˜7.7 associates to every confined boundaried -colorful graph a new one with the same -folio but with a bounded number of vertices. We denote this new confined boundaried -colorful graph by
Lemma 7.8.
There exists an algorithm that computes in time.
Proof.
The algorithm enumerates all confined boundaried -colorful graphs with and returns one with the same -topological-folio as By theorem˜7.7, such a graph exists. The computation of can be done in time, according to the main result of [42]. ∎
7.5 AMTs for colorful graphs
We now have all the necessary ingredients for our AMT’s. We conclude with the following corollary of theorem˜7.4 when applied for and taking into account corollary˜7.2, proposition˜7.5, along with lemma˜7.8 and theorem˜7.7.
Given a a functional problem is folio-representable if there exists a function such that is a representation of
Theorem 7.9.
Let and be a folio-representable, MSO-definable functional problem (via a formula and a ). There exists an algorithm that, given a -colorful graph computes in time, where is the maximum monodimensionality of a color in .
Proof.
Let be the -colorful graph such that for each vertex if and only if and is the color of maximum monodimensionality in Note that the monodimensionality of in is of order (provided that is a fixed number).
First, apply corollary˜7.2 to to obtain a star decomposition of with -width in time. Then apply theorem˜7.4: Condition (i) holds by proposition˜7.5 where and For Condition ii), we have that is a representation of and the algorithms computing and respectively are given by lemma˜7.8 and theorem˜7.7 for ∎
If now in the above proof we use corollary˜7.3 instead of corollary˜7.2, we replace by and we apply proposition˜7.6 instead of proposition˜7.5 we have the following.
Theorem 7.10.
Let and be a folio-representable, -definable functional problem (via formula and a ). There exists an algorithm that, given a -colorful graph computes in time, where is the maximum Hadwiger number of a color in .
Notice that theorem˜7.9 and theorem˜7.10 can be viewed as adaptations of proposition˜7.5 and proposition˜7.6 to optimization problems on colorful graphs, where the combinatorial condition captures how the colored vertices of the input should be distributed in rather than imposing a sparsity condition on the input graph that ignores the colors.
Folio representability for decision problems on colorful graphs.
A decision problem on -colorful graphs can be seen as a subset For completeness, we derive the notion of folio representability for decision problems from its functional counterpart as follows: A decision problem is folio representable if there exists a function such that for every -boundaried graph with boundary and every compatible confined -boundaried graph it holds that
where Recall that in the introduction we presented the decision problem analogues of theorem˜7.9 and theorem˜7.10, namely theorem˜1.8 and theorem˜1.9. These can be easily derived as follows: Let and let be a decision problem on -colorful graphs that is folio-representable and CMSO-definable (resp. -definable) by some formula We define a new formula on -colorful graphs so that if and only if the equivalence holds. Then, if and only if We then apply the algorithm of theorem˜7.9 (resp. theorem˜7.10) to compute the value of If this value is then while if it is then
8 Conclusion
As previously noted, the algorithmic contributions of this paper are limited to showcasing the algorithmic aspects of the colorful minor framework we developed. Below, we outline several promising directions for further research on problems involving colorful graphs.
8.1 Packing and covering colorful minors
As a general research project, one might consider the problem of computing the parameters and for different instantiations of (see section˜6.5 for the definitions). Next, we give examples of known problems that can be expressed by these parameters, given suitable choices of
Problems expressible via packing and covering colorful minors.
First of all, assume that is the -colorful rainbow and observe that by Menger’s theorem, for every -colorful graph it holds that and the two parameters are obviously computable in polynomial time. This is extended by lemma˜3.2 implying that, for every positive and are (linearly) equivalent when we consider the -colorful graph defined as follows: is the union of disjoint copies of and, for assigns to the vertex of the colors
Trying to push this further, one may consider a possible duality in the case of four sets and ask whether the following is correct: if there is no --linkage and no --linkage, both of order then there is either a --separator or a --separator of size Such a duality cannot exist, not even approximately, since it corresponds to the -colorful graph where assigns colors to one of the two vertices of and colors to the other one, and this -colorful graph is one of the obstructions for the Erdős Pósa property of -colorful graphs (the counterexample is given by figure˜11).
Let now be the -colorful rainbow Then is the maximum number of vertex disjoint paths between pairs of distinct vertices from This is the Maximum Disjoint T-Paths problem that can be solved in according to [92].
Notice that for the same asks whether there are vertices that, when removed, leave all remaining colored vertices in different connected components. This is the Node Multiway Cut with Deletable Terminals problem that is NP-complete even when the input graph is restricted to be planar [47].
We now proceed with an example of a problem that can be reduced to a covering problem on -colorful graphs.
This is Node Multicut with Deletable Terminals where the input is a graph a sequence of pairs of distinct terminals, and a The question is whether there is a set of at most vertices such that, for if then there is no - path in We now define a -colorful graph as follows. We first subdivide once all edges of and then add a set of new vertices and for and make the vertex adjacent to the vertices and We also define so that Let be a cycle on 4 vertices and let be a coloring assigning the color 1 to some of the vertices of It is easy to observe that is a yes instance of Node Multicut with Deletable Terminals if and only if
Constructing algorithms for packing and covering.
It is easy to observe that both parameters and are colorful minor-monotone, for every instantiation of Therefore, checking whether (resp. ) can be done by an algorithm running in time because of theorem˜1.7. However, as already commented after theorem˜1.7, this argument is not constructive and only yields the existence of the claimed algorithms.
One may construct an algorithm for checking whether as follows: We first check whether contains a -colorful rainbow using theorem˜1.6 in time. If this is the case, then we report a positive answer. If not, then we know that therefore, as is folio representable and -definable, we can apply theorem˜1.9 and construct an algorithm running deciding whether in time
On the other hand, a general constructive algorithm for checking whether requires more work. We may certainly begin by checking whether the -colorful rainbow is a colorful minor of and if so, we report negatively. Again, given that we may attempt to apply theorem˜1.9 for While is -definable, it is not folio representable, and therefore we cannot apply the replacements used in the proof of theorem˜7.4.
However, checking whether is still possible when For this, we may apply the replacement argument from theorem˜7.4 by using instead the stronger notion of a -folio of a boundaried graph We denote by - the collection of all -folios of the boundaried graphs derived from after removing vertices. As shown by Fomin, Lokshtanov, Panolan, Saurabh, and Zehavi in [32], one can construct an algorithm that, given outputs, in time, a new boundaried graph such that -- and the graph of excludes a clique as a minor, where the size of the excluded clique depends on a function of and (This follows by repeatedly applying the algorithm from [32, Theorem 6], which identifies irrelevant vertices.)
This enables a replacement argument that produces an equivalent instance whose graph has bounded Hadwiger number. The problem can then be solved in time, again using the decision version of proposition˜7.6.
To extend this argument to one would need to prove a “colorful” analogue of [32, Theorem 6], which would be able to identify a vertex that is irrelevant for every possible hitting set of size While one might reasonably believe such an extension is possible, the details of its construction appear to be quite intricate.
All algorithms discussed in this section have significant complexity, both in their parameter dependence and in their polynomial factors. Optimizing these algorithms – either in general or for specific, interesting instantiations of – could be a promising direction for future research.
8.2 Extending graph parameters to colorful ones
Recently, two main mechanisms have been proposed to extend treewidth to annotated graphs (i.e., -colorful graphs). These are the rainbow extension [84] and the torso extension [46, 45]. We propose here an extension of both mechanisms to -colorful graphs that applies to every minor-monotone parameter.
Let and let be a minor-monotone graph parameter. We define the torso -extension of and the rainbow -extension of as the -colorful graph parameters and , where
As mentioned in the introduction, , the torso treewidth of an annotated graph, has already been proposed—using different terminology—by Jansen and Swennenhuis in [46] and by Hodor, La, Micek, and Rambaud in [45]. Also, it is the the main combinatorial engine in the proof of theorem˜7.9.
On the other hand, , the rainbow treewidth of an annotated graph, is equivalent to the bidimensionality , which is , where denotes the maximum for which contains the -grid as a minor. This leads us to view (which, again, is equivalent to ) as the correct notion of bidimensionality for colorful graphs. In this sense, measures the “simultaneous presence” of all colors across the bidimensional surface of a grid.
Recall that, as already proved in [45], is equivalent to the maximum for which contains the -segregated grid as a colorful minor. Moreover, by theorem˜1.3, this equivalence extends to colorful graphs with any number of colors. That is, is equivalent to the maximum for which contains one of the parametrically-distinct -segregated grids as a colorful minor (see section˜6.3). This motivates us to interpret this quantity as the monodimensionality of a colorful graph, as it measures the “segregated presence” of all colors along the monodimensional boundary of a grid.
A natural question is when the two aforementioned mechanisms for generating colorful extensions yield equivalent parameters. Let be a minor-monotone graph parameter. It is easy to see that for every -colorful graph , we have , regardless of the choice of or . However, it is not generally the case that there exists a function such that for all . Thus, the question becomes:
For which minor-monotone parameters does it hold that and are equivalent?
We say that a graph parameter is unbounded in a class if the set is infinite. For two graph parameters and , we write if there is a function such that, for every graph , . We conjecture the following.
Conjecture 8.1.
Let and let be a minor-monotone parameter, where . Then
if , then and are equivalent if and only if is unbounded in outerplanar graphs;
if , then and are equivalent if and only if is unbounded in paths;
if , then and are not equivalent.
Proof of the “only if” directions of conjecture˜8.1.
We start with the case where . We use for the -segregated grid and we assume that is bounded on outerplanar graphs. Observe first that, for every , all rainbow minors of are outerplanar, therefore the value of on a is upper bounded by some constant that is independent of the value of . On the other side, we claim that the value of on is lower bounded by some (increasing) function of . To prove this claim, assume to the contrary, that there is some constant , such that for every , . This means that there is a where . But as , this means that is also always upper bounded by some constant . Notice now that is a -connected set of , i.e., and for all subsets with there are disjoint - paths in . Because , is a -connected set also in , which implies that is lower bounded by for some constant (see [20, Proposition 3]). This means that . We now have a contradiction by picking to be big enough so that and the claim holds. We just proved that the values of and may differ arbitrarily on . Therefore and we may conclude that and are not equivalent colorful graph parameters.
The proof for goes in a similar way. We use for the -segregated grid and we assume that . The fact that is bounded by a constant, no matter the choice of , follows by the fact that every rainbow minor of is the disjoint union of paths. Assume now that there is some , that does not depend on , and a set where (and therefore also ) and, for every component of , it holds that either or . Using this and the fact that there are vertex disjoint paths in between and , one can easily conclude that, provided that , for every component of , contains at most vertices from . This in turn implies that , where . The proof now goes as in the case where because of the fact that is a -connected set of which, for hight enough values of , contradicts the fact that .
For the case , notice that every rainbow parameter is bounded in any -segregated grid as these graphs do not contain rainbow minors. The fact that the torso extension of is unbounded on such graphs is a straightforward extension of the proof when , again using the fact that for every two distinct colors , there are vertex disjoint - paths. ∎
A proof of the “if” direction of conjecture˜8.1 (if exists) seems to be more involved and we leave it as an open problem.
A typical example where conjecture˜8.1 holds is pathwidth,131313We do not give the definitions of the parameters pathwidth and treedepth here. For our purposes, it is enough to mention that a graph class has bounded pathwidth/treedepth if and only if it does not contain all trees/paths. denoted by . While and are equivalent when (because of the results of [45]), this equivalence fails when because pathwidth is bounded in paths. For the case , notice that outperplanar graphs have unbounded pathwidth.
Another example is treedepth, denoted by . In this case, it is possible to verify that and are equivalent for both (for this one may apply theorem˜1.3 in order to bound torso treewidth and then recycle the proof of [67, Theorem 12.12]). For the cases , , notice that paths, as well as outperplanar graphs, have unbounded treedepth.
8.3 From faces to vortices
In this paper we presented some algorithmic consequences of two of our three structural theorems, i.e., theorem˜1.2 and theorem˜1.3. Clearly, the structure revealed by theorem˜1.4 is “intermediate” with respect to that of theorem˜1.2 and theorem˜1.3. It provides a structural decomposition of -colorful graphs of small bidimensionality, i.e., -colorful graphs excluding a large -colorful rainbow grid as a minor, which we believe should be useful in algorithmic design. As we already mentioned, the algorithmic consequences of theorem˜1.4 go beyond the objectives of this paper. However, we outline below some directions on the Steiner Tree problem that are inspired by the current research.
In the introduction, we commented on Steiner Tree as one of the most studied problems on annotated graphs. Using the terminology of this paper, Jansen and Swennenhuis in [46] proved that Steiner Tree, with input a -colorful graph can be solved in time where is the torso treewidth of Is torso treewidth the “best possible” structural parameter on that, when bounded, Steiner Tree can be solved in polynomial time? The answer is negative. This follows by the fact that a polynomial algorithm has been recently given by Groenland, Nederlof, and Koana in [41] for the case where excludes the -colorful rainbow as a colorful minor. As this -colorful graph belongs in the obstruction set because of corollary˜6.34, the results of [41] apply to all colorful minors of the -segregated grid on which torso treewidth may be arbitrarily large due to lemma˜5.4. We believe that the “limit” is given by the combinatorial bound of theorem˜1.4. We conjecture that the bidimensionality of in is the structural frontier for the complexity of Steiner Tree problem in classes that are colorful minor-closed. In particular we conjecture the following complexity dichotomy.
Conjecture 8.2.
Unless Steiner Tree can be solved in polynomial time in a -colorful minor-closed class if and only if has bounded bidimensionality.
Note that the negative part of the above conjecture already follows from the fact that the Steiner Tree problem is NP-complete in planar graphs [35]. We believe that the local structure provided by theorem˜1.4 should serve as a starting point for proving the positive part as well. In particular, we now have that all terminals in are confined to the apices and the “few” vortices of the near-embedding of Motivated by the fact that Steiner Tree can be solved in polynomial time when all terminals lie on “few” faces of a planar graph [22, 25, 55], the key challenge becomes how to generalize such face-based techniques to handle apices and vortices in near-embeddings of the kind provided by theorem˜1.4. We further anticipate that this methodological approach may lead to algorithmic dichotomies – such as the one conjectured above – for a broader class of problems on annotated graphs, where tractability arises from restricting terminals to few faces of the input graph.
Appendix A Some problems where our algorithmic results apply
In this section, we present a selection of natural optimization problems for which our algorithmic results apply. These problems are definable in CMSO and satisfy the conditions required by theorem˜7.9 or theorem˜7.10.
A.1 Problems where theorem˜7.9 can be applied
We now give some examples of optimization problems where theorem˜7.9 is applicable. For each of them, there is some formula in CMSO and a choice of such that if the input is the -colorful graph then answer is For all of them it is straightforward to verify that there exists a function such that is a representation of In particular then the question is about paths or cycles, then is enough, while when the problem is defined in terms of some graph then we may pick
We proceed with the list of problems. For each of them we may associate a decision problem by introducing an integer and asking whether the requested maximum/minimum number is at least/at most In each case, we give arguments why the corresponding decision version is NP-hard.
Spanning -Topological Minor
Input: A -colorful graph
Question: What is the maximum number of vertices of a subdivision of in may span?
Notice that when then the above problem yields Maximum Spanning Cycle that asks for the maximum number of vertices of that can be spanned by a cycle of The corresponding decision problem is NP-hard as, if we additionally set we obtain the Hamiltonian Cycle problem.
Longest Spanning Path/Cycle
Input: A -colorful graph
Question: What is the maximum path/cycle containing all vertices of ?
The corresponding decision problem is NP-hard as, if we additionally set we obtain the Longest Cycle/Longest Path problem.
Minimum/Maximum Number of Spanning Cycles
Input: A -colorful graph
Question: What is the minimum/maximum number of pairwise-disjoint cycles
spanning ?
The corresponding decision problem of the Minimum variant is NP-hard because it gives Hamiltonian Cycle when and The corresponding decision problem of the maximum variant reduces to the Disjoint Paths problem, in which we are given a graph along with a list of pairs of terminals and asks if contains pairwise vertex disjoint paths joining them. The reduction introduces a new vertex set to and makes each adjacent to and
Nailed Packing of -Subdivisions
Input: A -colorful graph
Question: what is the maximum
for which contains the disjoint union of
subdivisions of as a subgraph, each of which spans at least one vertex of ?
The corresponding decision problem, for is NP-hard because when this gives the Cycle Packing problem. This problem may be extended to -colorful graphs where we may add that each cycle spans all colors.
Alternating Spanning Cycle/Path Input: A -colorful graph Question: What is the maximum length cycle/path spanning such that no two vertices of the same set in appear consecutively along this path/cycle.
The corresponding decision problem is NP-hard because of the following reduction from Hamiltonian Cycle/Path: Given an instance of Hamiltonian Cycle/Path consider the instance of the decision version of Alternating Spanning Cycle/Path where is the bipartite double cover of (that is the tensor product of and ) and and are its two color classes.
We remark that the above problem can be extended for general -colorful graphs and for every pattern of color alternation.
Minimum Steiner -Forest Input: A -colorful graph Question: What is the minimum size of a subgraph of in which any two vertices belonging to the same set are connected by a path in ?
The corresponding decision problem, for the case is -hard, as proved in [87].
A.2 Problems where theorem˜7.10 can be applied
We now give two examples of optimization problems where theorem˜7.10 is applicable.
Unordered Linkability problems.
Given a graph a set and a graph on vertices, we say that is -linkable in if for every subset there is a and a collection of internally vertex-disjoint paths in between the pairs in that, when contracted to single edges, give a graph that is isomorphic to We consider the following optimization problem.
Unordered Linkability
Input: A -colorful graph and a graph on vertices.
Question: what is the biggest so that is -linkable in ?
Notice that the decision version of Unordered Linkability where is a cycle on vertices expresses (by asking whether ) is the Cyclability problem asking, given whether is -cyclable in that is every vertices of belong in some cycle of (see [21, 89, 43, 1, 31] for the combinatorial properties of -cyclable sets). In [36] it was proved that Cyclability, parameterized by is fixed parameter tractable (FPT) for planar graphs, while co-W[1]-hard in general. By FPT we mean that it admits an FPT-algorithm, i.e., an algorithm running in time.
Ordered Linkability problems.
We next define the “ordered” version of linkability. Let be a -colorful graph, and let be a graph with We say that is -linked in if, for every sequence of vertices in there is a and a collection of internally vertex-disjoint paths in between the pairs in that, when contracted to single edges, give a graph with vertex set that is isomorphic to via the isomorphism that maps to The combinatorics of -linked sets has been studied in [39, 57, 24, 40, 29]. In order to study the algorithmic properties of -linked sets, we consider the following optimization problem.
Ordered Linkability
Input: A -colorful graph and a graph on vertices.
Question: what is the biggest so that -linked in ?
In [37], a special case of Ordered Linkability problem was examined, namely the Disjoint Paths Linkability problem that is obtained if is the disjoint union of edges. It was proven in [37] that, when parameterized by this problem is not FPT, unless -linked sets for this particular where introduced by [90] in the late 60s and its graph-theoretical properties have been extensively studied in [60, 48, 79, 8, 23, 51, 86].
As observed in [37], the decision versions of Unordered Linkability and Ordered Linkability are -definable, therefore, from proposition˜7.6, they both admit an time algorithm. In other words, while no FPT-algorithms are expected when parameterized by they both become fixed parameter tractable (FPT), when they are further restricted to graphs of bounded Hadwiger number.
According to our results, towards achieving FPT-algorithms, we can do more than bounding the Hadwiger number of the input graph. For this, it can be easily seen that both these problems are folio-representable, therefore, from theorem˜1.9, we may deduce that they both admit an time algorithm. This reveals that it is not only the structure of that determines the parameterized complexity of these problems but the way the vertices in are distributed in
Packing and covering.
As commented in the section˜8, by using theorem˜7.10 we can construct algorithms running in time for the following decision problem (for every ).
Packing Minor Models of
Input: A -colorful graph and a
Question:
Does contain subgraphs each containing as a colorful minor?
Also, for as commented in section˜8.1, we have an algorithm running in time for the following decision problem.
Covering Minor Models of 1-colorful
Input: A -colorful graph and a
Question:
Does contain a set such that and
does not contain as a colorful minor?
We would like to conclude this section by commenting on how far the simple arguments we explained in section˜8.1 can go towards creating constructive algorithms for generalizations of Covering Minor Models of 1-colorful .
Let be any computable colorful minor-monotone 1-colorful graph parameter such that there is a constructive function where, for every 1-colorful graph it holds that Consider now the extension of Covering Minor Models of 1-colorful so that, instead of asking that we ask whether By using the same arguments as we explained in section˜8.1, one may construct an FPT-algorithm for this problem.
We emphasize that the collection of problems presented in this section is intended to be illustrative rather than exhaustive. Our aim is to showcase the diversity of settings – spanning connectivity, packing, spanning, and linkability – where our framework applies.